### v1

# Oscillations Make Neural Networks Robust to Quantization

Abstract
We challenge the prevailing view that oscillations in Quantization Aware Training (QAT) are merely undesirable artifacts caused by the Straight-Through Estimator (STE). Through theoretical analysis of QAT in linear models, we demonstrate that the gradient of the loss function can be decomposed into two terms: the original full-precision loss and a term that causes quantization oscillations. Based on these insights, we propose a novel regularization method that induces oscillations to improve quantization robustness. Contrary to traditional methods that focuses on minimizing the effects of oscillations, our approach leverages the beneficial aspects of weight oscillations to preserve model performance under quantization. Our empirical results on ResNet-18 and Tiny ViT demonstrate that this counter-intuitive strategy matches QAT accuracy at $\geq$ 3-bit weight quantization, while maintaining close to full precision accuracy at bits greater than the target bit. Our work therefore provides a new perspective on model preparation for quantization, particularly for finding weights that are robust to changes in the bit of the quantizer – an area where current methods struggle to match the accuracy of QAT at specific bits.  [^1] Source code available at: https://github.com/saintslab/OsciQuant
Section 1 Introduction
Quantization is the mapping of continuous values to discrete values. In neural networks, quantization reduces the computational complexity and memory requirements by representing weights and/or activations with fewer bits Gupta et al. (2015). In the case of weight only quantization, this means applying a quantizer $q(\cdot)$ to the network’s weights $\mathbf{w}$, with an additional implicit goal of maintaining the original performance i.e. $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$, where $\mathcal{L}(\cdot)$ is a loss function.
When training neural networks intended for quantization, an essential step during optimization is accounting for the effects of applying a quantizer to the weights. Quantization introduces a perturbation to the weights. For uniform quantizers, this is bounded by $\frac{s}{2}$, where $s$ is the scale factor. At higher bit widths ($\geq 8$ bits), this perturbation is small, and standard training procedures often yield weights that are resilient to quantization noise Nagel et al. (2021). In such cases, applying quantization after training, known as Post-Training Quantization (PTQ), is sufficient to maintain acceptable performance levels Nagel et al. (2021).
However, as we reduce the bit width to lower precision ($\leq 4$ bits), the quantization perturbation becomes more significant, and the model’s performance tends to degrade substantially after quantization. This is because the increased perturbation can lead to larger discrepancy between $q(\mathbf{w})$ and $\mathbf{w}$. To address this challenge, much research has gone into finding strategies to mitigate the effects of quantization on model accuracy, ensuring that the network remains accurate even after low-bit quantization.
Figure 1: Oscillatory behavior in Quantization-Aware Training (QAT) for a simple linear model. The figure shows a quantized linear model $f(x)=q(w)x$ with a single weight $w$, where $x=1$ and target output $y=0.75$. When doing squared loss with QAT an additional term is introduced to the gradient (Eq. 14), which causes w to oscillate around the quantization threshold. This oscillation results in $q(w)$ alternating between the 0 and 1 quantization bins.
Though many methods have been proposed for mitigating the accuracy degradation due to quantization, Quantization-Aware Training (QAT) Jacob et al. (2018) remains one of the most widely adopted approaches. QAT works by incorporating quantization effects directly into the training process - quantizing weights in the forward pass while using the Straight-Through Estimator (STE) Bengio et al. (2013) for gradient approximation during backpropagation. Research has identified an interesting phenomenon in QAT with the STE, known as weight oscillations, where the quantized weights alternate between two adjacent quantized states during training Défossez et al. (2021); Nagel et al. (2022). While traditionally viewed as a detrimental effect that should be suppressed through dampening or weight freezing techniques, there also exists evidence suggesting these oscillations might play a more nuanced role in the training dynamics of QAT.
We claim that weight oscillations during training are beneficial and that indeed they are the driving mechanism behind QAT. Our primary contributions that support this claim are:
  1. 1.
    we isolate the mechanism that leads to weight oscillations during QAT (Sec. 4);
  2. 2.
    we develop a regularization method that induces weight oscillations during training using this mechanism (Sec. 5);
  3. 3.
    we show experimentally that weight oscillations are sufficient for preserving performance after quantization on small-scale computer vision tasks (Sec. 6).
Since previous results have shown that weights oscillations are also necessary for good quantization performance with QAT (see Sec. 7 for details), and extrapolating from our experiments, our results suggest that weight oscillations capture all the beneficial effects of QAT while avoiding unintended side-effects. For instance, in our experiments our method avoids overfitting to the bit-width used during training, resulting in superior cross-quantization performance compared to QAT.
Section 2 Related Work
The most used strategy to minimize the impact of quantization on model accuracy is to minimize the quantization error. This can be achieved by adjusting the granularity of the quantizer—for instance, using per-channel Nagel et al. (2019) or block-wise quantization Dettmers et al. (2022) instead of per-tensor quantization. While these methods reduce quantization error without additional training, they come with increased storage requirements due to extra quantization parameters and may still fall short at very low bit widths, necessitating the combination with other approaches.
Consequently, extensive research has been dedicated to developing techniques that explicitly minimize the quantization error during optimization Hung et al. (2015); Hirose et al. (2017); Li et al. (2019); Choi et al. (2020); Han et al. (2021); Zhong et al. (2025). Given a model $\mathbf{w}$ the hope is that by ensuring $q(\mathbf{w})\approx\mathbf{w}$, we likely also have $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$, thereby preserving model accuracy after quantization.
An alternative and less explored approach involves training models to be robust to quantization perturbations without necessarily minimizing the quantization error itself. This means finding weights $\mathbf{w}$ such that $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$ even if $q(\mathbf{w})$ is not close to $\mathbf{w}$ Alizadeh et al. (2020); Chmiel et al. (2020). Such methods focus on enhancing the robustness of the model to the quantization error, leading to better performance at bits different than the ones used in the quantizer, which we will refer to as cross-bit quantization.
A third approach is to train supernets on the desired configurations of the quantizers Xu et al. ; Xu et al. (2023). This approach increases the training complexity and cost, which is not incurred by explicit regularization.
Despite these efforts, the aforementioned strategies often fall short of the accuracy obtained with QAT Jacob et al. (2018) at individual bits or indirectly rely upon QAT themselves. In short, QAT integrates the quantization process into the training loop allowing the model to adapt to the quantization effects directly. This is done by quantizing the weights during the forward pass and using techniques like the Straight-Through-Estimator (STE) to approximate the gradient of the quantizer (Which has a derivative of zero almost everywhere) during backpropagation Bengio et al. (2013).
Yet, there is limited understanding of how QAT affects model optimization and why it outperforms other methods. One phenomenon observed during QAT is weight oscillations Défossez et al. (2021); Nagel et al. (2022), which are periodic changes in the value of the quantized weight between two adjacent quantization levels. It is speculated in these works that that the abrupt changes in values caused by oscillations can interfere negatively with optimization. Oscillations are assumed to be undesirable side effects caused by the use of the STE during backpropagation, as the STE allows gradients to pass through the rounding operation in the quantizer, which has a gradient of zero almost everywhere Défossez et al. (2021); Nagel et al. (2022).
Several approaches have been suggested to mitigate oscillations, such as dampening or freezing the oscillating weights, which have shown improved accuracy Nagel et al. (2022); Gupta and Asthana (2024). However, the reported gains are sometimes marginal, and these methods may inadvertently also hinder the optimization process. For instance, Nagel et al. (2022) notes that freezing or dampening weights too early during training can hurt optimization, indicating that oscillations might contribute to finding better quantized minima. Liu et al. (2023) propose that weights with low oscillation frequency should be frozen, where as high-frequency ones should be left unfrozen, under the rational that high frequency means the network has little confidence in what value to quantize the weight to, where as low frequency means the optimal weight lies close to a quantization level.
Though QAT often provides the best accuracy for a given target bit, degradation to a lesser or greater extent exists when the bit of the quantizer is different to the one seen during training, ie. cross-bit quantization Alizadeh et al. (2020); Chmiel et al. (2020). This means QAT requires training and storing of weights for each desired bit width. This specialization can also pose challenges when deploying models across different hardware platforms, each potentially using different quantization schemes Reddi et al. (2020), making it difficult to develop models which can be easily quantized at deployment according to end-user requirements.
This makes robust quantization methods an interesting research avenue, especially if they could be improved to match the individual bit performance of QAT. In this work, we aim to deepen the understanding of how QAT influences model optimization, particularly focusing on the role of weight oscillations and their relation to robustness.
Section 3 Preliminaries
Section 3.1 Quantization
A quantizer divides a continuous input range into quantization bins, where each bin is represented by a specific quantization level. The boundaries between bins are called quantization thresholds. During quantization, any value within a bin is mapped to that bin’s quantization level. With a uniform quantizer, the step size (the distance between two adjacent quantization levels) is equal to the scale factor $s$.
We consider a uniform symmetric quantizer with a max-range scale factor. The quantization operation $q(\cdot)$ can then be expressed as
$q(\mathbf{w})$ $=s\cdot\left\lceil\frac{\mathbf{w}}{s}\right\rfloor$ (1)
[…]
The scale factor $s$ is set to cover the range of $\mathbf{w}$ as this removes the need for the usual clamping operation in the quantizer, while keeping the number of bins symmetric around 0:
$s$ $=\frac{\max(\lvert\alpha\rvert,\lvert\beta\rvert)}{2^{b-1}-1}$ (2)
Where $b$ is the bit in the quantizer and $\alpha,\beta$ are the min. and max. values respectively of the layer wise weight $\mathbf{w}$.
The quantization process introduces quantization error $\mathbf{\Delta}$, defined as the difference between the original and quantized values:
$\mathbf{\Delta}(\mathbf{w})$ $=\mathbf{w}-q(\mathbf{w})$ (3)
Due to the uniform quantizer, for all bins the absolute error is bounded between $20\leq|\mathbf{\Delta}|\leq s/$ 2, which is maximized at quantization thresholds and 0 at quantization levels.
Section 3.2 Quantization-Aware Training
While there exist many variants of QAT, fundamentally the forward pass is performed using the quantized weights $q(\mathbf{w})$ in most variants of QAT Jacob et al. (2018); Krishnamoorthi (2018), simulating the effect of using low-precision weights. In principle the gradient for the weights during QAT is given by:
$\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial\mathbf{w}}=\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial q(\mathbf{w})}\cdot\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}$ (4)
A problem with the above formulation is that the gradient of the rounding operation used in the quantizer is zero almost everywhere, causing the last term to interrupt gradient-based learning. A popular solution to this problem is to use the so-called Straight-Through Estimator (STE) Bengio et al. (2013). We define the STE to be the operator ^^$\frac{\hat{\partial}}{\hat{\partial}\mathbf{x}}$ such that ^$\frac{\hat{\partial}f}{\hat{\partial}x}$ is obtained by computing $\frac{\partial f}{\partial x}$ and in the resulting expression replacing $q^{\prime}$ (the derivative of $q$) by the constant function equal to $1$. In other words, if $\frac{\partial f}{\partial x}=g(\ldots,q^{\prime},\ldots)$ then ^$\frac{\hat{\partial}f}{\hat{\partial}x}=g(\ldots,1,\ldots)$.
Section 4 Oscillations in QAT
Previous studies have explored linear models to analyze the behavior of QAT and the phenomenon of weight oscillations Défossez et al. (2021); Nagel et al. (2022); Liu et al. (2023); Gupta and Asthana (2024). Inspired by these works, we analyze a linear regression model to gain theoretical insights into the optimization dynamics during QAT.
Consider a linear model with a single weight $w$, input $x$ and target $y$ $\in\mathbb{R}$. The quantized version of this model is defined as $f(x)=q(w)x$, where $q(\cdot)$ is the quantizer from Eq. 1. The quadratic loss for the quantized model is given by
$\mathcal{L}(q(w))=\frac{1}{2}(q(w)x-y)^{2}.$ (5)
Our goal in this section is to understand how QAT affects the full precision optimization process. For a given loss function $\mathcal{L}(\cdot)$ with quantized weights, we have
$\mathcal{L}(q(w))$ $=\mathcal{L}(w)+\mathcal{L}(q(w))-\mathcal{L}(w)$ (6)
We can then expand the difference caused by quantization as follows
$\delta_{\text{$\mathcal{L}$}}$ $=\mathcal{L}(q(w))-\mathcal{L}(w)$ (7)
$=\frac{1}{2}\left((q(w)x-y)^{2}-(wx-y)^{2}\right)$ (8)
$=\frac{1}{2}\left((q(w)x)^{2}-(wx)^{2}-2y(q(w)x-wx)\right)$ (9)
$=\frac{1}{2}\left(x^{2}\left(q(w)^{2}-w^{2}\right)\right)+\left(yx(w-q(w))\right)$ (10)
This expression decomposes the loss difference into a quadratic term $\frac{1}{2}x^{2}(q(w)^{2}-w^{2})$ and a linear term $yx(w-q(w))$.
Next we derive the gradient of $\delta_{\mathcal{L}}$ wrt. $w$:
$\frac{\partial\delta_{\mathcal{L}}}{\partial w}$ $=\frac{\partial}{\partial w}\bigg{(}\mathcal{L}(q(w))-\mathcal{L}(w)\bigg{)}$ (11)
$=\frac{\partial}{\partial w}\left(\frac{1}{2}x^{2}(q(w)^{2}-w^{2})+yx(w-q(w))\right)$ (12)
$=x^{2}\left(q(w)\frac{\partial q(w)}{\partial w}-w\right)+yx\left(1-\frac{\partial q(w)}{\partial w}\right)$ (13)
Using the STE and recalling that ^$\frac{\hat{\partial}q}{\hat{\partial}w}=1$ the expression of the STE gradient simplifies to
^$\frac{\hat{\partial}\delta_{\mathcal{L}}}{\hat{\partial}w}=x^{2}(q(w)-w)=-x^{2}\mathbf{\Delta}(w).$ (14)
To see how this gives rise to oscillations, for an arbitrary $w$, denote $w_{0}$ the upper discretization threshold $w_{0}=q(w)+s/2$. For $\varepsilon\in(0,s/2)$ note that we have $q(w_{0}-\varepsilon)=q(w)$ and $q(w_{0}+\varepsilon)=q(w)+s$ so that
$\mathbf{\Delta}(w_{0}+\varepsilon)$ $=q(w)+s/2+\varepsilon-(q(w)+s)$ (15)
$=-s/2+\varepsilon,$ (16)
$\mathbf{\Delta}(w_{0}-\varepsilon)$ $=q(w)+s/2-\varepsilon-q(w)$ (17)
$=s/2-\varepsilon.$ (18)
Assuming $x\neq 0$, the negative STE gradient “flips" from $-s/2$ to $s/2$ as the weight $w$ passes the quantization threshold $w_{0}$ from above, pushing the weight back towards the threshold. We note that the STE gradient is 00 at the special value $w=q(w)$, but the preceding argument shows that this is an unstable critical point and gradient noise will immediately cause the weights to move away from it. When combined with (stochastic) gradient descent and a finite discretization timestep we can identify this as the driving mechanism behind oscillations during training with QAT (Fig. 1).
We can also see how the dynamics lead to clustering around quantization thresholds by looking at the sign of $\mathbf{\Delta}$ for different values of $w$. For a weight $w$ let $d_{\text{low}}(w)$ and $d_{\text{up}}(w)$ denote the distance from $w$ to the upper and lower thresholds, $d_{\text{low}}(w)=w-\bigl{(}q(w)-\frac{s}{2}\bigr{)}=\mathbf{\Delta}(w)+\tfrac{s}{2}$ and $d_{\text{up}}(w)=\bigl{(}q(w)+\frac{s}{2}\bigr{)}-w=\tfrac{s}{2}-\mathbf{\Delta}(w)$ respectively. If $w$ is closest to the upper threshold we have
$d_{\text{up}}<d_{\text{low}}\Longrightarrow\tfrac{s}{2}-\mathbf{\Delta}<\mathbf{\Delta}+\tfrac{s}{2}\Longrightarrow\mathbf{\Delta}>0$ (19)
While if $w$ is closest to the lower threshold
$d_{\text{low}}<d_{\text{up}}\Longrightarrow\mathbf{\Delta}+\tfrac{s}{2}<\tfrac{s}{2}-\mathbf{\Delta}\Longrightarrow\mathbf{\Delta}<0$ (20)
We emphasize that this mechanism causes the weights to move towards the quantization thresholds (the edges of quantization "bins") as opposed to the quantization levels (the centers of the quantization "bins").
Section 5 Regularization Method
Based on our theoretical observations in the one weight linear model, we now investigate empirically if the mechanism in Eq. (14) is sufficient to introduce weight oscillations in neural networks.
From the quantization difference in Eq. 10 and the STE gradient derived in Eq. 14, we have:
$\frac{\partial\mathcal{L}(q(w))}{\partial w}$ $=\frac{\partial\mathcal{L}(w)}{\partial w}-x^{2}\mathbf{\Delta}(w)$ (21)
where the first term is the gradient of the original full-precision loss, and the second term causes the quantization oscillations in QAT.
In order to emulate the effects of QAT, we propose a regularization term so that the training objective becomes:
$\mathcal{L}(q(\mathbf{w}))=\mathcal{L}(\mathbf{w})+\mathcal{R_{\lambda}}(\mathbf{w})$ (22)
where we let the regularization term be similar to the quadratic term in Eq. (10):
$\mathcal{R_{\lambda}}(\mathbf{w})=\frac{\lambda}{2}\sum_{\ell}\frac{1}{n_{\ell}}\sum_{i=1}^{n_{\ell}}\left(q(w^{\ell}_{i})^{2}-(w^{\ell}_{i})^{2}\right).$ (23)
Here $\lambda\geq 0$ is a hyperparameter that controls the amount of regularization, $\ell$ ranges over the layers in the model and $i$ over the weights in each layer.
Using the STE, ^$\frac{\hat{\partial}q}{\partial\mathbf{w}}=1$, we have the following expression for the gradient:
^^$\frac{\hat{\partial}}{\hat{\partial}w^{\ell}_{i}}\mathcal{R_{\lambda}}(\mathbf{w})=\frac{\lambda}{n_{\ell}}\left(q(w^{\ell}_{i})-w^{\ell}_{i}\right)=-\frac{\lambda}{n_{\ell}}\mathbf{\Delta}(w_{i}^{\ell}).$ (24)
By the same reasoning as in Sec. 4 this pulls the weight $w_{i}^{\ell}$ towards the quantization threshold and causes the gradient to “flip" as $w_{i}^{\ell}$ crosses the threshold. We expect this to lead to oscillations based on the same mechanism as in the model from Sec. 4.
Figures 2 and  3 show the results of an experiment where we observe the weight distributions, and measured the oscillations, during training of a neural network (ResNet-18) with varying degrees of regularization, respectively. For comparison purposes the figures also shows the weight distributions and oscillations observed during training with QAT. Using the definition of an oscillation established in Nagel et al. (2022), we count an oscillation at epoch $i>1$ if $q(w_{t})\neq q(w_{t-1})$ and the direction of the change in the quantized space is opposite to that of the previous change. We note though that this method of counting misses the first threshold crossing during an oscillation A.5
Our first observation is that QAT displays more oscillations – also seen as clustering around the quantization threshold in Fig. 2-a) – than a baseline model without QAT or regularization (corresponding to $\lambda=0$ in Fig. 3-b)) . As we increase $\lambda$ we observe that the number of oscillations as well as the clustering increases. This confirms that our regularizer can indeed induce oscillations similar to QAT during the training of deep neural networks. At $\lambda=1$ (Fig. 3-c)) the number of oscillations observed with our regularizer is similar to the behaviour of QAT, lending support to our hypothesis that the mechanism in (15) is indeed at the root of the oscillations observed when training neural networks with QAT.
Figure 2: Weight distribution analysis of ResNet-18’s first convolutional layer after 50 epochs of training from scratch. a) Weight distribution under QAT with a 3-bit quantizer. b)-d) Our proposed regularization approach with a 3-bit quantizer at varying regularization strengths ($\lambda=0,1,10$, from left to right). When $\lambda=0$, the training reduces to standard optimization. The QAT distribution (leftmost) exhibits the characteristic threshold clustering behavior. As $\lambda$ increases, we observe progressively stronger clustering of weights around quantization thresholds, illustrating the relationship between regularization strength and weight clustering.
Figure 3: The plots show the distribution of weights with oscillation counts $>0$ when training with a) QAT and b)-d) our regularizer for different values of $\lambda$. Here $\lambda=0$ corresponds to a full precision model where our regularizer has no influence on training. The y-axis represents the percentage of total weights in the first convolutional layer of a ResNet-18 trained from scratch for 50 epochs, while the x-axis shows the oscillation count. Following the oscillation definition from Nagel et al. (2022), we count oscillations at each epoch during training. The results demonstrate that QAT produces a significantly higher proportion of oscillating weights compared to $\lambda=0$. Furthermore, we observe that as we increase $\lambda$ a greater percentage of weights oscillates.
Section 6 Experiments & Results
In this section we empirically try to answer the question: is it sufficient to induce weight oscillations during training in order to get the benefits of QAT?
We answer this question mostly affirmatively for ResNet and Vision Transformer architectures, based on the results of training ResNet-18 and Tiny ViT on the CIFAR-10 dataset. This is both in a training-from-scratch setting and when fine-tuning pretrained models. In all our experiments we use the regularizer $\mathcal{R}_{\lambda}$ defined in Eq. (23) to induce oscillations.
In the following subsections we first describe the experimental setup, then we present the accuracy results from training-from-scratch and fine-tuning models trained with different quantization levels for the quantizer in $\mathcal{R}_{\lambda}$ or QAT and finally, we present the cross-bit accuracy of the fine-tuned models. We train models at ternary (3 possible values: -1, 0, 1), 3-bit and 4-bit. This is in line with contemporary research, where the emphasis lies on quantization at 4-bit and below since the challenges of maintaining accuracy are more significant compared to quantization at higher bit widths.
Section 6.1 Experimental setup
We conducted our experiments using the CIFAR-10 dataset Krizhevsky et al. (2009) without data augmentation. We evaluated three architectures; A multi-layer perceptron with 5 hidden layers and 256 neurons per layer (MLP5), ResNet-18 He et al. (2016) and Tiny Vision transformer (Tiny ViT) Wu et al. (2022).
For each architecture we used the Adam optimizer Kingma (2014) and tested multiple configurations: A baseline model to establish optimal floating-point accuracy and post-training quantization (PTQ) performance, a model with QAT and a model with our approach. The two latter configurations are trained using a ternary, 3-bit, and 4-bit quantizer.
Training from Scratch For the MLP5 architecture, we used a learning rate of $10^{-3}$ and regularization parameter $\lambda$=1. The ResNet-18 was trained with a learning rate of $10^{-3}$ and $\lambda$=0.75 (see Appx. A.2 for our hyperparameter selection). We modified the ResNet-18 architecture by replacing the input layer with a smaller $3\times 3$ kernel and adapting the final layer for 10-class classification of both ResNet-18 and Tiny ViT. Training proceeded for a maximum of 100 epochs with early stopping triggered after 10 epochs without improvement in validation performance. For quantized models, we monitored the quantized validation accuracy at the target bit precision, while for the baseline, we tracked floating-point accuracy.
Fine-tuning We fine-tuned two ImageNet-1k Deng et al. (2009) pre-trained models on CIFAR-10: a Tiny ViT (learning rate: $10^{-4}$, $\lambda$=1) and a ResNet-18 (learning rate: $10^{-3}$, $\lambda$=1). To maintain compatibility with the pre-trained architectures, we upsampled CIFAR-10 images to $224\times 224$ pixels. The $\lambda$ parameter selection process for Tiny ViT is detailed in Appx. A.2. Fine-tuning continued for up to 200 epochs, with early stopping after 30 epochs without improvement, using the same accuracy metrics as training from scratch.
Quantization We implemented weight quantization using a per-tensor uniform symmetric quantizer as defined in Eq. 1. The quantization range was determined by computing minimum and maximum values per layer. In our implementation of ResNet-18 (11M parameters) all layers except batch normalization were quantized, covering 99.96% of parameters. For Tiny ViT (5.5M parameters) quantization was applied to MLP, Self-Attention, and key-query-value projection layers, encompassing 97.18% of parameters. And lastly for the MLP5 model all layers were quantized.
Section 6.2 Training-from-scratch
Table 1 shows the results from training an MLP and ResNet-18 from scratch on the CIFAR-10 dataset. Our regularization method (OsciQuant) demonstrates improvements compared to the PTQ baseline from ternary quantization. More importantly, it also matches the performance of QAT at bit widths of 3 and 4.
For both models we see that at 3-bit and 4-bit, our method exhibits similar performance as QAT but with less variability, while not differing significantly in the average number of training epochs required. With both models, QAT and OsciQuant are competitive with the full-precision baseline, although we observe an increased number of training epochs. Notably, both OsciQuant and QAT significantly outperform PTQ when applied to the full precision baseline.
Model Quantization method Accuracy Mean Epochs
MLP5 Baseline FP32 51.43 $\pm$ 0.39 14
Ternary PTQ 10.00 $\pm$ 0.02 14
Ternary QAT 49.20 $\pm$ 1.34 24
Ternary OsciQuant 36.49 $\pm$ 0.51 14
3-bit PTQ 20.97 $\pm$ 5.64 14
3-bit QAT 50.53 $\pm$ 1.43 33
3-bit OsciQuant 48.48 $\pm$ 0.29 15
4-bit PTQ 46.50 $\pm$ 0.76 14
4-bit QAT 51.39 $\pm$ 0.60 26
4-bit OsciQuant 50.72 $\pm$ 0.47 19
ResNet-18 Baseline FP32 83.26 $\pm$ 1.07 24
Ternary PTQ 10.00 $\pm$ 0.01 24
Ternary QAT 79.62 $\pm$ 6.42 42
Ternary OsciQuant 61.5 $\pm$ 1.82 56
3-bit PTQ 77.79 $\pm$ 4.0 24
3-bit QAT 82.51 $\pm$ 2.14 37
3-bit OsciQuant 81.77 $\pm$ 0.46 41
4-bit PTQ 82.11 $\pm$ 1.21 24
4-bit QAT 82.66 $\pm$ 2.57 28
4-bit OsciQuant 83.74 $\pm$ 0.59 32
Table 1: Comparison of accuracy when training from scratch on CIFAR-10. Results show classification accuracy and mean training epochs for MLP5 and ResNet-18 across different quantization approaches and bit-widths. Results is means and standard deviations over 5 random seeds.
Model Quantization method Accuracy Mean Epochs
ResNet-18 Baseline FP32 88.50 $\pm$ 0.64 4
Ternary PTQ 10.01 $\pm$ 0.01 4
Ternary QAT 77.02 $\pm$ 7.57 47
Ternary OsciQuant 44.59 $\pm$ 3.30 35
3-bit PTQ 10.28 $\pm$ 0.48 4
3-bit QAT 85.69 $\pm$ 1.83 25
3-bit OsciQuant 84.94 $\pm$ 1.59 27
4-bit PTQ 35.56 $\pm$ 9.05 4
4-bit QAT 87.71 $\pm$ 1.14 26
4-bit OsciQuant 87.08 $\pm$ 0.72 24
Tiny ViT Baseline FP32 96.11 $\pm$ 0.31 6
Ternary PTQ 9.39 $\pm$ 1.11 6
Ternary QAT 73.53 $\pm$ 0.77 140
Ternary OsciQuant 13.51 $\pm$ 1.32 28
3-bit PTQ 11.56 $\pm$ 1.99 6
3-bit QAT 88.13 $\pm$ 0.60 131
3-bit OsciQuant 88.68 $\pm$ 1.08 108
4-bit PTQ 21.57 $\pm$ 5.33 6
4-bit QAT 94.96 $\pm$ 0.33 57
4-bit OsciQuant 94.82 $\pm$ 0.51 90
Table 2: Comparison of accuracy when fine-tuning on models pre-trained on ImageNet-1k. Results show classification accuracy and mean training epochs for MLP5 and ResNet-18 across different quantization approaches and bit-widths. Results is means and standard deviations over 5 random seeds.
Section 6.3 Fine-tuning
Table 2 summarizes the test accuracies for fine-tuning using our OsciQuant method and QAT on ResNet-18 and Tiny ViT architectures. The observations are roughly in line with the results observed for training from scratch in the previous section with the exception of the number of epochs required for fine-tuning.
On the ResNet architecture both QAT and our model train for significantly longer than the full precision baseline. As is the case for training from scratch, we see an increase in ternary performance compared to PTQ, but QAT still ourperforms our method in the ternary setting. Our regularization and QAT show comparable performance when quantized at 3 bits and 4 bits, while achieving test accuracy close to the full precision model at 4-bits.
The general trend regarding accuracy is identical for the vision transformer experiments, while we again note the high number of epochs require for both methods when fine-tuning, compared to the full precision baseline.
Section 6.4 Robustness to cross-bit quantization
As described above, the goal of our proposed regularization term is to train a model that maintains performance after quantization. Since the regularization term involves a quantization operator, we need to choose the quantization level in the regularization term. In this experiment we evaluated the robustness of our method and QAT towards quantization at levels different from the ones used during training.
For OsciQuant, we applied a regularization term with the training bit width during training and applied PTQ after training finished at a different quantization level. For QAT we trained using the training bit width and afterwards applied PTQ to the latent weights. For each method we also evaluated the corresponding model without PTQ, directly using the latent weights for inference (reported as FP32).
Table 3 shows the results from the experiment. A first observation is that the models produced by our method consistently achieve nearly full-precision accuracy when quantized at 8-bit or when used without quantization, irrespective of the quantization level used during training. This contrasts with QAT, which produces a viable 8-bit or full-precision model only when trained with at least 4-bit.
Furthermore we see that our method mostly maintains performance when trained at 3 or 4-bit and quantized at bit level of 3 or 4-bit. QAT also achieves this for Tiny ViT but for ResNet, the accuracy of QAT trained at 3-bit and quantized at other bit widths is barely above random guessing.
Regarding training with ternary quantization, we see that our method produces models that achieve near full precision performance for ResNet when quantized at 3-bit or higher. Ternary training for ViT is somewhat peculiar in that it fails to produce a model that is viable when quantized to ternary, whereas the performance of the resulting models starts to show a high level of variability at 4-bit and finally reaches close to full-precision accuracy at 8-bit. In contrast, for both ResNet and ViT, the performance of QAT degrades completely to random guessing when trained with ternary quantization and evaluated at any other quantization level.
Model Train bit $\downarrow$ / Eval. bit $\rightarrow$ FP32 Ternary 3-bit 4-bit 8-bit
ResNet-18 Baseline (PTQ) 88.50 $\pm$ 0.64 10.01 $\pm$ 0.01 10.28 $\pm$ 0.48 35.56 $\pm$ 9.05 88.45 $\pm$ 0.64
Ternary QAT 10.39 $\pm$ 0.71 77.02 $\pm$ 7.57 9.75 $\pm$ 0.77 10.03 $\pm$ 0.51 10.35 $\pm$ 0.63
Ternary OsciQuant 87.44 $\pm$ 0.56 44.59 $\pm$ 3.30 85.42 $\pm$ 1.13 87.03 $\pm$ 0.65 87.42 $\pm$ 0.56
3-bit QAT 16.89 $\pm$ 4.97 10.01 $\pm$ 0.04 85.69 $\pm$ 1.83 17.42 $\pm$ 4.96 16.56 $\pm$ 4.32
3-bit OsciQuant 87.86 $\pm$ 0.42 20.19 $\pm$ 10.74 84.94 $\pm$ 1.59 87.56 $\pm$ 0.38 87.86 $\pm$ 0.42
4-bit QAT 87.75 $\pm$ 1.13 10.13 $\pm$ 0.29 82.08 $\pm$ 6.25 87.71 $\pm$ 1.14 87.76 $\pm$ 1.12
4-bit OsciQuant 87.85 $\pm$ 0.49 11.91 $\pm$ 0.87 85.57 $\pm$ 1.10 87.08 $\pm$ 0.72 87.87 $\pm$ 0.49
Tiny ViT Baseline (PTQ) 96.11 $\pm$ 0.31 9.39 $\pm$ 1.11 11.56 $\pm$ 1.99 21.57 $\pm$ 5.33 96.03 $\pm$ 0.34
Ternary QAT 10.62 $\pm$ 1.29 73.53 $\pm$ 0.77 11.52 $\pm$ 1.82 11.13 $\pm$ 1.75 10.61 $\pm$ 1.26
Ternary OsciQuant 95.79 $\pm$ 0.58 13.51 $\pm$ 1.32 12.53 $\pm$ 3.66 54.93 $\pm$ 27.32 95.76 $\pm$ 0.59
3-bit QAT 86.94 $\pm$ 0.91 19.78 $\pm$ 6.04 88.13 $\pm$ 0.60 86.69 $\pm$ 0.62 86.95 $\pm$ 0.89
3-bit OsciQuant 96.47 $\pm$ 0.11 9.48 $\pm$ 1.64 88.68 $\pm$ 1.08 95.35 $\pm$ 0.18 96.50 $\pm$ 0.11
4-bit QAT 95.14 $\pm$ 0.29 $1$$\pm$ 1.84 59.86 $\pm$ 19.95 94.96 $\pm$ 0.33 95.13 $\pm$ 0.28
4-bit OsciQuant 96.54 $\pm$ 0.09 11.90 $\pm$ 1.29 70.23 $\pm$ 12.75 94.82 $\pm$ 0.51 96.55 $\pm$ 0.09
Table 3: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on CIFAR-10. Grey background is the target-bit accuracy. Models are trained using different quantization methods (QAT and ours) and bit-widths (ternary, 3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32. The grey diagonal shows the results for the bit used during training. Results are means and standard deviations over 5 random seeds. All significant differences between QAT and OsciQuant are shown in bold face.
Section 7 Discussion
We have shown that training with weight oscillations induced via regularization is sufficient in most cases to maintain performance after quantization for ResNet and Tiny ViT. This begs the question whether weight oscillations are also a necessary part of the QAT training process. Indeed, some previous work already points towards this. There are examples claiming that both dampening and/or freezing of oscillations too early in the training process is detrimental to performance after quantization Nagel et al. (2022); Han et al. (2021). And in other case presented in Liu et al. (2023), freezing only the low frequency oscillating weights improves performance. This suggests that weight oscillations are both a necessary and sufficient part of QAT, at least in the early phases of the training process. This further supports our hypothesis that oscillations in QAT have a positive effect on quantization robustness.
Additionally, there might be further benefits to our regularization approach compared to QAT. Our method aims to isolate this crucial part of the training process. This is arguably a more principled approach compared to QAT, where quantization during training combined with STE can lead to a number of side-effects beyond oscillations, which can be highly non-intuitive. We present a simple example in the Appendix Sec. A.1 where replacing a single scalar weight by a product of two scalar weights leads to a non-trivial change in training dynamics when using QAT with the STE.
On the other hand, while it is not clear what the additional effects are during QAT, we do note two consistent deviations from the QAT performance when using our regularization method: QAT outperforms regularization at ternary quantization, whereas our regularization method outperforms QAT in cross-bit accuracy for the ternary and 3-bit case. In A.4, we see how it seems that the cross-bit performance for QAT is upper-bounded by the target-bit performance, which might explain the subpar QAT performance at cross-bit compared to our regularization method which seems bounded by the full precision accuracy. Additionally we can note that while it is stated in Alizadeh et al. (2020); Chmiel et al. (2020) that QAT is not robust to cross-bit quantization, A.4 shows that for some cases the robustness is tied closely to how long the model is trained after the target bit accuracy has converged.
Finally we note in A.2 that in the ResNet-18 model, we see similar results for the hyperparameter sweep for different $\lambda$s, which might suggest that the key for robustness is the presence of oscillations and not their precise nature.
Limitations In our experiments we observed that the robustness to cross-bit quantization improves in later training epochs. In order to further improve robustness one might consider an early stopping criterion that evaluates the performance on cross-bit quantization, which was not done in this work. The same approach could also increase cross-bit quantization robustness of QAT although to a lesser degree than for our method.
We performed our experiments on the CIFAR-10 dataset which might make it more difficult to compare our results with other published works that provide benchmark results for other datasets such as ImageNet-1k.
Section 8 Conclusion
Based on the analysis of a toy model we proposed the hypothesis that weight oscillations during training in deep neural networks make the model robust to quantization.
In Sections 4 and 5 we explain on a toy model how training with QAT and STE leads to oscillations and propose a regularizer that encourages this oscillating behaviour. We confirm that as we increase the strength of the regularization, we empirically observe the appearance of clustering together with oscillations.
Finally we experimentally confirm that the regularizer indeed leads to consistent robustness towards quantization for quantization levels above ternary. Our regularization method achieves comparable performance to QAT above ternary quantization when quantizing to the target-bit seen during optimizing and shows increased robustness compared to QAT in cross-bit quantization with bits greater than the target-bit used in the quantizer during training. All this being evidence of our hypothesis.
Our insights on weight oscillations and their role in quantization robustness open new horizons for model quantization approaches. Our regularization method especially creates interesting possibilities for cross-bit robustness, potentially making our regularization method more appealing than QAT when the goal is to deploy or relase a single set of weights that works across different bit widths or maybe even quantizers. While the regularizer used in our experiments should be viewed as an initial step, we expect that quantization robustness could be further improved by developing oscillation-inducing methods that are adaptive to different learning rates, layer statistics or phases of the training process.
Acknowledgments
Authors thank Tong Chen, Jákup Svøðstein and Sebastian Hammer Eliassen for useful discussions.
JW, BP and RS are partly funded by European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284, No. 101070408 and No. 101189771. RS also acknowledges funding received under Independent Research Fund Denmark (DFF) under grant 4307-00143B.
Appendix A Appendix
Section A.1 2-layer with single weights
Figure 4: We repeat the toy model experiments, but this time with 2 weights, taking into account that the linear term is no longer 0 in the gradient. We notice at epoch 15 and 18 where the prediction of the quantized model is greater than y, the effect of the terms flip for $w_{2}$.
Consider a linear model $f(x)=w_{2}w_{1}x$, with $w_{1},w_{2}$, input $x$, and target $y\in\mathbb{R}$. The quantized version of this model is defined as $f_{q}(x)=q(w_{2})q(w_{1})x$, where $q(\cdot)$ is the quantizer from Eq. 1. The quadratic loss for the model is given by
$\mathcal{L}(f(x))=\frac{1}{2}\bigl{(}w_{2}w_{1}x-y\bigr{)}^{2}$
The difference compared to full-precision optimization is then given as
$\delta_{\mathcal{L}}$ $=\mathcal{L}(f_{q}(x))-\mathcal{L}(f(x))$ (25)
$=\frac{1}{2}\Bigl{[}\bigl{(}q(w_{2})q(w_{1})x-y\bigr{)}^{2}-\bigl{(}w_{2}w_{1}x-y\bigr{)}^{2}\Bigr{]}$ (26)
$=\frac{1}{2}\Bigl{[}\bigl{(}q(w_{2})q(w_{1})x\bigr{)}^{2}-\bigl{(}w_{2}w_{1}x\bigr{)}^{2}-2y\bigl{(}q(w_{2})q(w_{1})x-w_{2}w_{1}x\bigr{)}\Bigr{]}$ (27)
$=\frac{1}{2}x^{2}\Bigl{[}q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr{]}+yx\Bigl{[}w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr{]}$ (28)
The loss difference decomposes into:
$\underbrace{\frac{1}{2}x^{2}\Bigl{(}q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr{)}}_{\text{quadratic term}}\quad+\quad\underbrace{yx\Bigl{(}w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr{)}}_{\text{linear term}}$
Taking the derivative of $\mathcal{L}$ with respect to $w_{1}$:
$\frac{\partial\delta_{\mathcal{L}}}{\partial w_{1}}$ $=\frac{\partial}{\partial w_{1}}\Bigl{(}\mathcal{L}(f_{q}(x))-\mathcal{L}(f(x))\Bigr{)}$ (29)
$=\frac{\partial}{\partial w_{1}}\biggl{[}\frac{1}{2}x^{2}\Bigl{(}q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr{)}+yx\Bigl{(}w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr{)}\biggr{]}$ (30)
$=x^{2}\Bigl{[}q(w_{2})^{2}q(w_{1})\frac{\partial q(w_{1})}{\partial w_{1}}-w_{2}^{2}w_{1}\Bigr{]}+yx\Bigl{[}w_{2}-q(w_{2})\frac{\partial q(w_{1})}{\partial w_{1}}\Bigr{]}$ (31)
Using the STE approximation from Eq. 4, we get:
$\frac{\partial\hat{\delta}_{\mathcal{L}}}{\partial w_{1}}=x^{2}\Bigl{[}q(w_{2})^{2}q(w_{1})-w_{2}^{2}w_{1}\Bigr{]}+yx\Bigl{[}w_{2}-q(w_{2})\Bigr{]}$ (32)
We note that the linear term is no longer zero in the gradient and thus for a model consisting of 2 single weight layers we see that there is additional effects from QAT other than oscillations. Additionally because of the non-linearity of the rounding operation, even with the absence of a non-linear activation function, we can no longer reduce the model to a single weight.
Section A.2 Hyperparameters
A.2.1 ResNet-18
$\lambda$ 3-bit (%) Ternary (%)
0.25 68.77 $\pm$ 0.19 47.85 $\pm$ 5.51
0.50 69.47 $\pm$ 1.11 46.77 $\pm$ 4.83
0.75 70.08 $\pm$ 0.40 46.86 $\pm$ 3.01
1.00 66.20 $\pm$ 4.05 47.33 $\pm$ 2.06
1.25 69.31 $\pm$ 0.32 43.14 $\pm$ 6.62
1.50 68.96 $\pm$ 0.30 46.73 $\pm$ 3.91
1.75 69.92 $\pm$ 0.11 47.02 $\pm$ 4.19
Figure 5: Mean over 3 runs of the best validation accuracy for different lambdas. Training a ResNet-18 from scratch. Both ternary and 3-bit is at $10^{-3}$ learning rate and 50% of the data used for training. The plot shows three learning rates, where we for each learning ratue evaluate with the $\lambda$s in the rhs. table. The colored background covers the range between the maximum and minimum value of the quantized validation accuracy with the given $\lambda$s.
In Fig. 5 we see the results of a hyperparameter search over different learning rates and $\lambda$s for a ResNet-18 model. There is a clear trend of seeing the best performance at a learning rate of $10^{-3}$. We note that interestingly there is a comparable performance for a wide range of $\lambda$s, indicating that it is the presence of oscillations which is important for quantization robustness, and not the exact frequency of oscillations.
A.2.2 Tiny ViT
$\lambda$ 3-bit (%) Ternary (%)
0.01 18.85 -
0.5 85.21 15.10
0.75 87.68 -
1.0 90.29 13.04
2.0 89.31 14.16
2.5 - 13.70
5.0 - 14.20
Figure 6: Validation accuracy at different $\lambda$ values and the corresponding best validation accuracies for 3-bit and 2-bit configurations for a single run. Learning rate is set to 1e-4 for fine-tuning. For the 2-bit we test higher $\lambda$ but still see no improvemenet in accuracy. We note how all the $\lambda$s lies close to each other, except for the low of $10^{-2}$
Fig. 6 We note how also the Tiny Vit seems to allow for a wide range of $\lambda$s even though we this time note that $\lambda=1$ performs significantly better than the others.
Section A.3 Epochs and cross-bit robustness
Figure 7: Left is the validation accuracy during training of a ViT with QAT at different bits, right is for our regularization. Both QAT and regularization is trained with a 3-bit quantizer. We note how the order of convergences for cross-bit changes between QAT and our model and that QAT cross-bit robustness especially depends on number of epochs trained.
There is an interesting interaction between number of epochs trained and robustness both of our method and QAT. We note how QAT converges first for the target-bit and then over time also converges for the 4 and 8-bit. Additionally we see that QAT seems upper-bounded by the target-bit performance, while this is not the case for our metho. Fig. 7,
Section A.4 Convergence behaviour of Tiny ViT
Figure 8: Regularization with a 3-bit quantizer on a Tiny ViT. We note the peculiar behaviour of the orange line, which is the validation accuracy on the target-bit performance. The performances cycles between $\approx 90\%$ and $10\%$, while the full precision accuracy (The model evaluated without quantized weights) stays some-what stable.
Fig. 8 shows the convergence behaviour of the full precision weights and the quantized weights at target-bit. We note how the Tiny ViT displays a peculiar convergence behaviour, where the accuracy will break, only to go up again. In the Tiny Vit model we quantize the self-attention layers, it is already noted in  Liu et al. [2023] that ViTs are especially vulnerable to quantization of the query and key of a self-attention layer, which might be related to the convergence behaviour we see.
Section A.5 Counting oscillations
Figure 9: The plots shows the weights with a total oscillation count $>0$ in a baseline model, a QAT model and a model regularized with $\lambda=1$ respectively. The y-axis is the percentage of the total weights in the first convolutional layer in a ResNet-18 trained from scratch for 50 epochs. For the baseline model we log the full precision weights at each epoch and then apply the quantizer after training, where as for the regularized model we simply log both the full precision and quantized weights at each epoch.
Using the counting method proposed by Nagel et al. [2022], we notice the biased odd distribution of the bins. In the plot used for the main paper, we count each bin the histogram in iterations of 2. Additionally we note that the proposed method misses the first oscillation count in each weight history.

### v2

@@
 Abstract
-We challenge the prevailing view that oscillations in Quantization Aware Training (QAT) are merely undesirable artifacts caused by the Straight-Through Estimator (STE).
-Through theoretical analysis of QAT in linear models, we demonstrate that the gradient of the loss function can be decomposed into two terms: the original full-precision loss and a term that causes quantization oscillations.
-Based on these insights, we propose a novel regularization method that induces oscillations to improve quantization robustness.
-Contrary to traditional methods that focuses on minimizing the effects of oscillations, our approach leverages the beneficial aspects of weight oscillations to preserve model performance under quantization.
-Our empirical results on ResNet-18 and Tiny ViT demonstrate that this counter-intuitive strategy matches QAT accuracy at $\geq$ 3-bit weight quantization, while maintaining close to full precision accuracy at bits greater than the target bit.
-Our work therefore provides a new perspective on model preparation for quantization, particularly for finding weights that are robust to changes in the bit of the quantizer – an area where current methods struggle to match the accuracy of QAT at specific bits.
-[^1] Source code available at: https://github.com/saintslab/OsciQuant
-Section 1 Introduction
-Quantization is the mapping of continuous values to discrete values.
-In neural networks, quantization reduces the computational complexity and memory requirements by representing weights and/or activations with fewer bits Gupta et al.
-(2015).
-In the case of weight only quantization, this means applying a quantizer $q(\cdot)$ to the network’s weights $\mathbf{w}$, with an additional implicit goal of maintaining the original performance i.e.
-$\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$, where $\mathcal{L}(\cdot)$ is a loss function.
-When training neural networks intended for quantization, an essential step during optimization is accounting for the effects of applying a quantizer to the weights.
-Quantization introduces a perturbation to the weights.
-For uniform quantizers, this is bounded by $\frac{s}{2}$, where $s$ is the scale factor.
-At higher bit widths ($\geq 8$ bits), this perturbation is small, and standard training procedures often yield weights that are resilient to quantization noise Nagel et al.
-(2021).
-In such cases, applying quantization after training, known as Post-Training Quantization (PTQ), is sufficient to maintain acceptable performance levels Nagel et al.
-(2021).
-However, as we reduce the bit width to lower precision ($\leq 4$ bits), the quantization perturbation becomes more significant, and the model’s performance tends to degrade substantially after quantization.
-This is because the increased perturbation can lead to larger discrepancy between $q(\mathbf{w})$ and $\mathbf{w}$.
-To address this challenge, much research has gone into finding strategies to mitigate the effects of quantization on model accuracy, ensuring that the network remains accurate even after low-bit quantization.
-Figure 1: Oscillatory behavior in Quantization-Aware Training (QAT) for a simple linear model.
-The figure shows a quantized linear model $f(x)=q(w)x$ with a single weight $w$, where $x=1$ and target output $y=0.75$.
-When doing squared loss with QAT an additional term is introduced to the gradient (Eq.
-14), which causes w to oscillate around the quantization threshold.
-This oscillation results in $q(w)$ alternating between the 0 and 1 quantization bins.
-Though many methods have been proposed for mitigating the accuracy degradation due to quantization, Quantization-Aware Training (QAT) Jacob et al.
-(2018) remains one of the most widely adopted approaches.
-QAT works by incorporating quantization effects directly into the training process - quantizing weights in the forward pass while using the Straight-Through Estimator (STE) Bengio et al.
-(2013) for gradient approximation during backpropagation.
-Research has identified an interesting phenomenon in QAT with the STE, known as weight oscillations, where the quantized weights alternate between two adjacent quantized states during training Défossez et al.
-(2021); Nagel et al.
-(2022).
-While traditionally viewed as a detrimental effect that should be suppressed through dampening or weight freezing techniques, there also exists evidence suggesting these oscillations might play a more nuanced role in the training dynamics of QAT.
-We claim that weight oscillations during training are beneficial and that indeed they are the driving mechanism behind QAT. Our primary contributions that support this claim are:
+We challenge the prevailing view that weight oscillations observed during Quantization Aware Training (QAT) are merely undesirable side-effects and argue instead that they are an essential part of QAT.
+We show in a univariate linear model that QAT results in an additional loss term that causes oscillations by pushing weights away from their nearest quantization level.
+Based on the mechanism from the analysis, we then derive a regularizer that induces oscillations in the weights of neural networks during training.
+Our empirical results on ResNet-18 and Tiny Vision Transformer, evaluated on CIFAR-10 and Tiny ImageNet datasets, demonstrate across a range of quantization levels that training with oscillations followed by post-training quantization (PTQ) is sufficient to recover the performance of QAT in most cases.
+With this work we provide further insight into the dynamics of QAT and contribute a novel insight into explaining the role of oscillations in QAT which until now have been considered to have a primarily negative effect on quantization.
+[^1] Source code is available at https://github.com/saintslab/osc_reg.
+1 Introduction
+Deep neural networks have grown increasingly successful at solving difficult modelling tasks, but are also increasingly becoming more expensive to use.
+As model sizes balloon into the hundreds of millions of parameters, the cost of inference has become a significant bottleneck, particularly for deployment on edge devices or usage in large-scale services sevilla2022compute.
+To reduce this cost, quantization of model weights has emerged as a prominent strategy nagel2021white.
+Weight quantization, however, comes with its own cost, namely a drop in accuracy.
+Reducing precision introduces quantization error, and naive approaches like Post-Training Quantization (PTQ) often lead to sharp degradations in model performance at low bit widths.
+To combat this degradation in accuracy, many methods have been proposed, usually centered around minimizing the quantization error BridgeDeepLearning; minimizeQuantError; APTQ; Choi_2020; ImprovingLowBit; zhong2024mbquantnovelmultibranchtopology.
+Figure 1: Oscillatory behavior during QAT in a one weight linear model $\hat{y}=q(w)x$, with weight $w$, quantized weight $q(w)$, input $x=1$ and target $y=0.75$.
+The gradient of the loss term of this model during QAT can be decomposed into two terms: $\nabla L(w)$ and $\nabla\delta_{L}=q(w)-w$, where the latter term is what differentiates QAT from just optimizing the full-precision loss $L(w)$.
+During QAT, $\nabla L(w)$ always points towards $y$, while $\nabla\delta_{L}$ introduces a dynamic which pushes $w$ towards the nearest bin threshold.
+This causes $w$ to oscillate when $y$ is not exactly on a quantization level.
+In the above case this makes $q(w)$ alternate between 0 and 1.
+Note the frequency of oscillations of $q(w)$ lets the quantized weight on average to converge to 0.75.
+Yet these approaches fall short of Quantization-Aware Training (QAT) jacob2017quantization; krishnamoorthi2018quantizing, which integrates quantization directly into the training loop, yielding models that remain robust even under low-precision constraints.
+However, despite its empirical success, the underlying behavior of QAT remains poorly understood.
+In particular, QAT often exhibits oscillations in the weights during optimization, while regular training does not.
+Rather than settling into a stable state, quantized weights fluctuate between adjacent quantization levels.
+These oscillations are widely seen as undesirable artifacts.
+Several works have tried to suppress these weight oscillations mainly through dampening or weight freezing pseudoQuantNoise; nagel2022overcoming; huang2023quantization; gupta2023reducing; vitoscillations.
+This paper poses the question: Have we misunderstood the nature of oscillations in QAT?
+To the best of our knowledge, we are the first to argue that weight oscillations in QAT are not a flaw, but rather an essential feature of the training dynamics.
+Using a toy model, we show that oscillations arise from a gradient component in QAT which pushes weights towards the nearest quantization threshold – this stands in contrast to aligning weights at the quantization level, which would minimize the quantization error.
+We then induce such oscillations to push the weights towards the nearest thresholds deliberately during the model training.
+This is realized as a simple regularization term that pushes weights towards their nearest quantization threshold.
+Surprisingly, this leads to models that are robust to quantization and in most cases we find that these regularization-induced oscillations recover the accuracy obtained with QAT.
+Additionally, in many cases, oscillations outperform QAT under cross-bit evaluation (i.e.
+testing at precision levels not simulated during training).
+These findings further deepen our understanding of QAT and point to a more nuanced role for weight oscillations; suggesting they may be beneficial to QAT rather than being solely detrimental as argued in most existing literature.
+Our primary contributions supporting this claim are:
   1. 1.
-    we isolate the mechanism that leads to weight oscillations during QAT (Sec. 4);
+    Using a univariate model we illustrate how the Straight Through Estimator (STE) leads to oscillations and clustering of model weights during QAT (Sec. 3);
   2. 2.
-    we develop a regularization method that induces weight oscillations during training using this mechanism (Sec. 5);
+    We show experimentally that by using a mechanism inspired by the toy model, we can induce oscillations and clustering during training in neural networks (Sec. 4).
   3. 3.
-    we show experimentally that weight oscillations are sufficient for preserving performance after quantization on small-scale computer vision tasks (Sec. 6).
-Since previous results have shown that weights oscillations are also necessary for good quantization performance with QAT (see Sec.
-7 for details), and extrapolating from our experiments, our results suggest that weight oscillations capture all the beneficial effects of QAT while avoiding unintended side-effects.
-For instance, in our experiments our method avoids overfitting to the bit-width used during training, resulting in superior cross-quantization performance compared to QAT.
-Section 2 Related Work
-The most used strategy to minimize the impact of quantization on model accuracy is to minimize the quantization error.
-This can be achieved by adjusting the granularity of the quantizer—for instance, using per-channel Nagel et al.
-(2019) or block-wise quantization Dettmers et al.
-(2022) instead of per-tensor quantization.
-While these methods reduce quantization error without additional training, they come with increased storage requirements due to extra quantization parameters and may still fall short at very low bit widths, necessitating the combination with other approaches.
-Consequently, extensive research has been dedicated to developing techniques that explicitly minimize the quantization error during optimization Hung et al.
-(2015); Hirose et al.
-(2017); Li et al.
-(2019); Choi et al.
-(2020); Han et al.
-(2021); Zhong et al.
-(2025).
-Given a model $\mathbf{w}$ the hope is that by ensuring $q(\mathbf{w})\approx\mathbf{w}$, we likely also have $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$, thereby preserving model accuracy after quantization.
-An alternative and less explored approach involves training models to be robust to quantization perturbations without necessarily minimizing the quantization error itself.
-This means finding weights $\mathbf{w}$ such that $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$ even if $q(\mathbf{w})$ is not close to $\mathbf{w}$ Alizadeh et al.
-(2020); Chmiel et al.
-(2020).
-Such methods focus on enhancing the robustness of the model to the quantization error, leading to better performance at bits different than the ones used in the quantizer, which we will refer to as cross-bit quantization.
-A third approach is to train supernets on the desired configurations of the quantizers Xu et al.
-; Xu et al.
-(2023).
-This approach increases the training complexity and cost, which is not incurred by explicit regularization.
-Despite these efforts, the aforementioned strategies often fall short of the accuracy obtained with QAT Jacob et al.
-(2018) at individual bits or indirectly rely upon QAT themselves.
-In short, QAT integrates the quantization process into the training loop allowing the model to adapt to the quantization effects directly.
-This is done by quantizing the weights during the forward pass and using techniques like the Straight-Through-Estimator (STE) to approximate the gradient of the quantizer (Which has a derivative of zero almost everywhere) during backpropagation Bengio et al.
-(2013).
-Yet, there is limited understanding of how QAT affects model optimization and why it outperforms other methods.
-One phenomenon observed during QAT is weight oscillations Défossez et al.
-(2021); Nagel et al.
-(2022), which are periodic changes in the value of the quantized weight between two adjacent quantization levels.
-It is speculated in these works that that the abrupt changes in values caused by oscillations can interfere negatively with optimization.
-Oscillations are assumed to be undesirable side effects caused by the use of the STE during backpropagation, as the STE allows gradients to pass through the rounding operation in the quantizer, which has a gradient of zero almost everywhere Défossez et al.
-(2021); Nagel et al.
-(2022).
-Several approaches have been suggested to mitigate oscillations, such as dampening or freezing the oscillating weights, which have shown improved accuracy Nagel et al.
-(2022); Gupta and Asthana (2024).
-However, the reported gains are sometimes marginal, and these methods may inadvertently also hinder the optimization process.
-For instance, Nagel et al.
-(2022) notes that freezing or dampening weights too early during training can hurt optimization, indicating that oscillations might contribute to finding better quantized minima.
-Liu et al.
-(2023) propose that weights with low oscillation frequency should be frozen, where as high-frequency ones should be left unfrozen, under the rational that high frequency means the network has little confidence in what value to quantize the weight to, where as low frequency means the optimal weight lies close to a quantization level.
-Though QAT often provides the best accuracy for a given target bit, degradation to a lesser or greater extent exists when the bit of the quantizer is different to the one seen during training, ie.
-cross-bit quantization Alizadeh et al.
-(2020); Chmiel et al.
-(2020).
-This means QAT requires training and storing of weights for each desired bit width.
-This specialization can also pose challenges when deploying models across different hardware platforms, each potentially using different quantization schemes Reddi et al.
-(2020), making it difficult to develop models which can be easily quantized at deployment according to end-user requirements.
-This makes robust quantization methods an interesting research avenue, especially if they could be improved to match the individual bit performance of QAT.
-In this work, we aim to deepen the understanding of how QAT influences model optimization, particularly focusing on the role of weight oscillations and their relation to robustness.
-Section 3 Preliminaries
-Section 3.1 Quantization
+    We empirically confirm using the CIFAR-10 krizhevsky2009learning and Tiny-ImageNet Le2015TinyIV datasets on a multi-layer perceptron (MLP), ResNet-18 he2016deep and Tiny Vision transformer (Tiny ViT) wu2022tinyvit that introducing oscillations through regularization in most cases recovers the accuracy of QAT (Sec.
+5).
+2 Preliminaries and Related Work
+Quantization:
 A quantizer divides a continuous input range into quantization bins, where each bin is represented by a specific quantization level.
@@
 With a uniform quantizer, the step size (the distance between two adjacent quantization levels) is equal to the scale factor $s$.
-We consider a uniform symmetric quantizer with a max-range scale factor. The quantization operation $q(\cdot)$ can then be expressed as
-$q(\mathbf{w})$ $=s\cdot\left\lceil\frac{\mathbf{w}}{s}\right\rfloor$ (1)
-[…]
-The scale factor $s$ is set to cover the range of $\mathbf{w}$ as this removes the need for the usual clamping operation in the quantizer, while keeping the number of bins symmetric around 0:
-$s$ $=\frac{\max(\lvert\alpha\rvert,\lvert\beta\rvert)}{2^{b-1}-1}$ (2)
-Where $b$ is the bit in the quantizer and $\alpha,\beta$ are the min. and max. values respectively of the layer wise weight $\mathbf{w}$.
+We consider a uniform symmetric, scalar quantizer, $q(\cdot)$, that can then be expressed as:
+$q(\mathbf{w})$ $=s\cdot\left\lceil\frac{\mathbf{w}}{s}\right\rfloor.$ (1)
+Here, $s$ represents the scale factor and $\left\lceil\cdot\right\rfloor$ denotes the rounding operation.
+The scale factor $s$ is set to cover the range of $\mathbf{w}$ as this removes the need for the clamping operation $\text{clamp}(q(\textbf{w});\alpha,\beta)$ in the quantizer.
+The function clamp(.) restricts $\alpha\leq q(w_{i})\leq\beta$ and any values less or greater than is set to $\alpha$ and $\beta$ respectively.
+Additionally we set the number of positive and negative quantization levels to be $2^{b-1}-1$, so we have symmetric number of levels around 0:
+$s$ $=\frac{\max(\lvert\textbf{w}\rvert)}{2^{b-1}-1},$ (2)
+where $b$ is the number of bits the quantizer can use.
 The quantization process introduces quantization error $\mathbf{\Delta}$, defined as the difference between the original and quantized values:
-$\mathbf{\Delta}(\mathbf{w})$ $=\mathbf{w}-q(\mathbf{w})$ (3)
-Due to the uniform quantizer, for all bins the absolute error is bounded between $20\leq|\mathbf{\Delta}|\leq s/$ 2, which is maximized at quantization thresholds and 0 at quantization levels.
-Section 3.2 Quantization-Aware Training
-While there exist many variants of QAT, fundamentally the forward pass is performed using the quantized weights $q(\mathbf{w})$ in most variants of QAT Jacob et al.
-(2018); Krishnamoorthi (2018), simulating the effect of using low-precision weights.
-In principle the gradient for the weights during QAT is given by:
-$\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial\mathbf{w}}=\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial q(\mathbf{w})}\cdot\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}$ (4)
-A problem with the above formulation is that the gradient of the rounding operation used in the quantizer is zero almost everywhere, causing the last term to interrupt gradient-based learning.
-A popular solution to this problem is to use the so-called Straight-Through Estimator (STE) Bengio et al.
+$\mathbf{\Delta}(\mathbf{w})$ $=\mathbf{w}-q(\mathbf{w}).$ (3)
+Due to the uniform quantizer each component of the absolute error (due to quantizing $w_{i}$) is bounded between $0\leq|\mathbf{\Delta}_{i}|\leq s/2$, for all the bins.
+This is maximized at quantization thresholds and is 0 at quantization levels.
+Straight-Through Estimator (STE):
+STE is a heuristic gradient approximation technique commonly used to enable backpropagation through functions $f(\cdot)$ whose gradients are zero almost everywhere bengio2013ste.
+In the forward pass, the function $f(\mathbf{x})$ is applied as usual.
+During the backward pass, however, its gradient is approximated by replacing the derivative of $f$ with the constant $1$.
+Formally, STE defines the approximate gradient of a function $f$ as
+$\frac{{\partial}f}{{\partial}\textbf{x}}\approx\frac{\hat{\partial}f}{\hat{\partial}\textbf{x}}=1.$ (4)
+Quantization-Aware Training (QAT):
+In QAT, the network maintains full-precision weights for learning but quantizes them during the forward pass before computation.
+This allows training to proceed with gradients applied to full-precision weights, while the network effectively operates on quantized values.
+While there exist many variants of QAT, the typical forward pass is performed using the quantized weights $q(\mathbf{w})$ jacob2017quantization; krishnamoorthi2018quantizing.
+The gradient for the weights during QAT is given by
+$\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial\mathbf{w}}=\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial q(\mathbf{w})}\cdot\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}.$ (5)
+A problem with the above formulation is that the gradient of the quantizer $\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}$ is zero almost everywhere, causing the last term to interrupt gradient-based learning.
+A widely used solution to this problem is to use the STE (Eq.
+(4)).
+Weight Oscillations:
+We use the definition from nagel2022overcoming which defines a weight oscillation during QAT to occur in iteration $t$ if it satisfies both the following conditions:
+  1. 1.
+    The quantized value of the weight needs to change between iterations i.e. $q(w_{t})\neq q(w_{t-1})$.
+  2. 2.
+    The direction of the change in the quantized domain needs to be the opposite than that of its previous change i.e.
+$\mathrm{sign}(\Gamma_{t})\neq\mathrm{sign}(\Gamma_{\tau})$, where $\tau$ is the iteration of the last change, and $\Gamma_{t}=q(w_{t})-q(w_{t-1})$ is the direction of the change.
+Related Work:
+Minimizing the quantization error is the most commonly used strategy to reduce the impact of quantization on model accuracy.
+Extensive research has been dedicated to developing techniques that explicitly minimize the quantization error during optimization BridgeDeepLearning; minimizeQuantError; APTQ; Choi_2020; ImprovingLowBit; zhong2024mbquantnovelmultibranchtopology.
+Despite these efforts, the aforementioned strategies often fall short of the accuracy obtained with QAT jacob2017quantization at individual bits or indirectly rely upon QAT themselves.
+However, there is limited understanding of how QAT affects model optimization and why it outperforms other methods.
+One phenomenon observed during QAT is weight oscillations pseudoQuantNoise; nagel2022overcoming, which are periodic changes in the value of the quantized weight between two adjacent quantization levels.
+It is speculated in these works that the abrupt changes in values caused by oscillations can interfere negatively with optimization.
+Oscillations are assumed to be undesirable side effects caused by the use of the STE during backpropagation, as the STE allows gradients to pass through the rounding operation in the quantizer, which has a gradient of zero almost everywhere pseudoQuantNoise; nagel2022overcoming.
+Several approaches have been suggested to mitigate oscillations by either freezing or dampening pseudoQuantNoise; nagel2022overcoming; huang2023quantization; gupta2023reducing; vitoscillations.
+However, the reported accuracy gains are sometimes marginal, and these methods may inadvertently also hinder the optimization process.
+For instance, nagel2022overcoming notes that freezing or dampening weights too early during training can hurt optimization, indicating that oscillations might contribute to finding better quantized minima of the loss.
+vitoscillations propose that weights with low oscillation frequency should be frozen, where as high-frequency ones should be left unfrozen, under the rationale that high frequency means the network has little confidence in what value to quantize the weight to, whereas low frequency means the optimal weight lies close to a quantization level.
+3 Oscillations in QAT
+Previous studies have explored linear models to analyze the behavior of QAT and the phenomenon of weight oscillations pseudoQuantNoise; nagel2022overcoming; vitoscillations; gupta2023reducing.
+Inspired by these works, we also analyze a linear regression model to gain insights into the optimization dynamics during QAT.
+3.1 Toy Model
+Consider a linear model with a single weight $w$, input $x$ and target $y$ $\in\mathbb{R}$.
+The quantized version of this model is defined as $\hat{y}=q(w)x$, where $q(\cdot)$ is the quantizer from Eq.
 (1).
-We define the STE to be the operator ^^$\frac{\hat{\partial}}{\hat{\partial}\mathbf{x}}$ such that ^$\frac{\hat{\partial}f}{\hat{\partial}x}$ is obtained by computing $\frac{\partial f}{\partial x}$ and in the resulting expression replacing $q^{\prime}$ (the derivative of $q$) by the constant function equal to $1$.
-In other words, if $\frac{\partial f}{\partial x}=g(\ldots,q^{\prime},\ldots)$ then ^$\frac{\hat{\partial}f}{\hat{\partial}x}=g(\ldots,1,\ldots)$.
-Section 4 Oscillations in QAT
-Previous studies have explored linear models to analyze the behavior of QAT and the phenomenon of weight oscillations Défossez et al.
-(2021); Nagel et al.
-(2022); Liu et al.
-(2023); Gupta and Asthana (2024).
-Inspired by these works, we analyze a linear regression model to gain theoretical insights into the optimization dynamics during QAT.
-Consider a linear model with a single weight $w$, input $x$ and target $y$ $\in\mathbb{R}$.
-The quantized version of this model is defined as $f(x)=q(w)x$, where $q(\cdot)$ is the quantizer from Eq.
-1.
 The quadratic loss for the quantized model is given by
-$\mathcal{L}(q(w))=\frac{1}{2}(q(w)x-y)^{2}.$ (5)
-Our goal in this section is to understand how QAT affects the full precision optimization process. For a given loss function $\mathcal{L}(\cdot)$ with quantized weights, we have
-$\mathcal{L}(q(w))$ $=\mathcal{L}(w)+\mathcal{L}(q(w))-\mathcal{L}(w)$ (6)
-We can then expand the difference caused by quantization as follows
-$\delta_{\text{$\mathcal{L}$}}$ $=\mathcal{L}(q(w))-\mathcal{L}(w)$ (7)
-$=\frac{1}{2}\left((q(w)x-y)^{2}-(wx-y)^{2}\right)$ (8)
-$=\frac{1}{2}\left((q(w)x)^{2}-(wx)^{2}-2y(q(w)x-wx)\right)$ (9)
-$=\frac{1}{2}\left(x^{2}\left(q(w)^{2}-w^{2}\right)\right)+\left(yx(w-q(w))\right)$ (10)
-This expression decomposes the loss difference into a quadratic term $\frac{1}{2}x^{2}(q(w)^{2}-w^{2})$ and a linear term $yx(w-q(w))$.
-Next we derive the gradient of $\delta_{\mathcal{L}}$ wrt. $w$:
-$\frac{\partial\delta_{\mathcal{L}}}{\partial w}$ $=\frac{\partial}{\partial w}\bigg{(}\mathcal{L}(q(w))-\mathcal{L}(w)\bigg{)}$ (11)
-$=\frac{\partial}{\partial w}\left(\frac{1}{2}x^{2}(q(w)^{2}-w^{2})+yx(w-q(w))\right)$ (12)
-$=x^{2}\left(q(w)\frac{\partial q(w)}{\partial w}-w\right)+yx\left(1-\frac{\partial q(w)}{\partial w}\right)$ (13)
-Using the STE and recalling that ^$\frac{\hat{\partial}q}{\hat{\partial}w}=1$ the expression of the STE gradient simplifies to
-^$\frac{\hat{\partial}\delta_{\mathcal{L}}}{\hat{\partial}w}=x^{2}(q(w)-w)=-x^{2}\mathbf{\Delta}(w).$ (14)
-To see how this gives rise to oscillations, for an arbitrary $w$, denote $w_{0}$ the upper discretization threshold $w_{0}=q(w)+s/2$.
-For $\varepsilon\in(0,s/2)$ note that we have $q(w_{0}-\varepsilon)=q(w)$ and $q(w_{0}+\varepsilon)=q(w)+s$ so that
-$\mathbf{\Delta}(w_{0}+\varepsilon)$ $=q(w)+s/2+\varepsilon-(q(w)+s)$ (15)
-$=-s/2+\varepsilon,$ (16)
-$\mathbf{\Delta}(w_{0}-\varepsilon)$ $=q(w)+s/2-\varepsilon-q(w)$ (17)
-$=s/2-\varepsilon.$ (18)
-Assuming $x\neq 0$, the negative STE gradient “flips" from $-s/2$ to $s/2$ as the weight $w$ passes the quantization threshold $w_{0}$ from above, pushing the weight back towards the threshold.
-We note that the STE gradient is 00 at the special value $w=q(w)$, but the preceding argument shows that this is an unstable critical point and gradient noise will immediately cause the weights to move away from it.
+$\mathcal{L}(q(w))=\frac{1}{2}(\hat{y}-y)^{2}=\frac{1}{2}(q(w)x-y)^{2}.$ (6)
+Our goal in this section is to understand how QAT affects the full-precision optimization process. For a given loss function $\mathcal{L}(\cdot)$ with quantized weights, we have
+$\mathcal{L}(q(w))$ $=\mathcal{L}(w)+\mathcal{L}(q(w))-\mathcal{L}(w).$ (7)
+We can then expand the difference in loss caused by quantization as follows
+$\delta_{\text{$\mathcal{L}$}}$ $=\mathcal{L}(q(w))-\mathcal{L}(w)=\frac{1}{2}\left((q(w)x-y)^{2}-(wx-y)^{2}\right)$ (8)
+$=\frac{1}{2}\left(x^{2}\left(q(w)^{2}-w^{2}\right)\right)+\left(yx(w-q(w))\right).$ (9)
+This expression decomposes the loss difference $\delta_{\text{$\mathcal{L}$}}$ into a quadratic term $\frac{1}{2}x^{2}(q(w)^{2}-w^{2})$ and a linear term yx(wq(w))yx(w-q(w)).
+Next we derive the gradient of $\delta_{\mathcal{L}}$ wrt. $w$
+$\frac{\partial\delta_{\mathcal{L}}}{\partial w}$ $=\frac{\partial}{\partial w}\bigg(\mathcal{L}(q(w))-\mathcal{L}(w)\bigg)=\frac{\partial}{\partial w}\left(\frac{1}{2}x^{2}(q(w)^{2}-w^{2})+yx(w-q(w))\right)$ (10)
+$=x^{2}\left(q(w)\frac{\partial q(w)}{\partial w}-w\right)+yx\left(1-\frac{\partial q(w)}{\partial w}\right).$ (11)
+Using the STE and recalling that $\frac{\hat{\partial}q}{\hat{\partial}w}=1$ the expression of the STE gradient simplifies to [^2] Note that there is no clamping in the quantizer because of the scale factor Eq.
+(2).
+$\frac{\hat{\partial}\delta_{\mathcal{L}}}{\hat{\partial}w}=x^{2}(q(w)-w)=-x^{2}\mathbf{\Delta}(w).$ (12)
+3.2 Oscillation Mechanism
+To see how the observations in Sec.
+3.1 give rise to oscillations, let $w_{0}$ denote the upper quantization threshold for an arbitrary weight $w$, defined as $w_{0}=q(w)+s/2$.
+For $\varepsilon\in(0,s/2)$, note that we have $q(w_{0}-\varepsilon)=q(w)$ and $q(w_{0}+\varepsilon)=q(w)+s$, so that
+$\mathbf{\Delta}(w_{0}+\varepsilon)$ $=q(w)+s/2+\varepsilon-(q(w)+s)=-s/2+\varepsilon,$ (13)
+$\mathbf{\Delta}(w_{0}-\varepsilon)$ $=q(w)+s/2-\varepsilon-q(w)=s/2-\varepsilon.$ (14)
+Assuming $x\neq 0$, the negative STE gradient “flips” from s/2-s/2 to +s/2+s/2 as the weight $w$ passes the quantization threshold $w_{0}$, pushing the weight back towards the threshold.
+We note that the STE gradient is zero at the special value $w=q(w)$, but the preceding argument shows that this is an unstable critical point and gradient noise will immediately cause the weights to move away from it.
 When combined with (stochastic) gradient descent and a finite discretization timestep we can identify this as the driving mechanism behind oscillations during training with QAT (Fig.
 1).
-We can also see how the dynamics lead to clustering around quantization thresholds by looking at the sign of $\mathbf{\Delta}$ for different values of $w$.
+3.3 Weight Clustering
+We can also see how the dynamics described above can lead to weight clustering around quantization thresholds by looking at the sign of $\mathbf{\Delta}$ for different values of $w$.
 For a weight $w$ let $d_{\text{low}}(w)$ and $d_{\text{up}}(w)$ denote the distance from $w$ to the upper and lower thresholds, $d_{\text{low}}(w)=w-\bigl(q(w)-\frac{s}{2}\bigr)=\mathbf{\Delta}(w)+\tfrac{s}{2}$ and $d_{\text{up}}(w)=\bigl(q(w)+\frac{s}{2}\bigr)-w=\tfrac{s}{2}-\mathbf{\Delta}(w)$ respectively.
 If $w$ is closest to the upper threshold we have
-$d_{\text{up}}<d_{\text{low}}\Longrightarrow\tfrac{s}{2}-\mathbf{\Delta}<\mathbf{\Delta}+\tfrac{s}{2}\Longrightarrow\mathbf{\Delta}>0$ (19)
+$d_{\text{up}}<d_{\text{low}}\Longrightarrow\tfrac{s}{2}-\mathbf{\Delta}<\mathbf{\Delta}+\tfrac{s}{2}\Longrightarrow\mathbf{\Delta}>0.$ (15)
 While if $w$ is closest to the lower threshold
-$d_{\text{low}}<d_{\text{up}}\Longrightarrow\mathbf{\Delta}+\tfrac{s}{2}<\tfrac{s}{2}-\mathbf{\Delta}\Longrightarrow\mathbf{\Delta}<0$ (20)
-We emphasize that this mechanism causes the weights to move towards the quantization thresholds (the edges of quantization "bins") as opposed to the quantization levels (the centers of the quantization "bins").
-Section 5 Regularization Method
-Based on our theoretical observations in the one weight linear model, we now investigate empirically if the mechanism in Eq. (14) is sufficient to introduce weight oscillations in neural networks.
-From the quantization difference in Eq. 10 and the STE gradient derived in Eq. 14, we have:
-$\frac{\partial\mathcal{L}(q(w))}{\partial w}$ $=\frac{\partial\mathcal{L}(w)}{\partial w}-x^{2}\mathbf{\Delta}(w)$ (21)
+$d_{\text{low}}<d_{\text{up}}\Longrightarrow\mathbf{\Delta}+\tfrac{s}{2}<\tfrac{s}{2}-\mathbf{\Delta}\Longrightarrow\mathbf{\Delta}<0.$ (16)
+We emphasize that this mechanism causes the weight to move towards the quantization thresholds (the edges of quantization “bins”) as opposed to the quantization levels (the centers of the quantization “bins”).
+As shown above, the magnitude of the pull towards the threshold increases as the weight approaches it.
+The weight will therefore eventually cross the threshold and start oscillating, unless $L(w)$ and $\delta_{L}$ exactly cancel out, which is unlikely to happen with a finite gradient descent step size.
+4 Regularization Method
+Our theoretical observations in the linear model in Sec.
+3, show that the oscillation component is the only part that differentiates QAT from standard full-precision training.
+We now confirm empirically that the mechanism in Eq.
+(12) is sufficient to introduce weight oscillations during training of neural networks, and study if this also results in QAT-like behavior with respect to the quantization noise.
+From the quantization difference in Eq. (9) and the STE gradient derived in Eq. (12), we have
+$\frac{\partial\mathcal{L}(q(w))}{\partial w}$ $=\frac{\partial\mathcal{L}(w)}{\partial w}-x^{2}\mathbf{\Delta}(w),$ (17)
 where the first term is the gradient of the original full-precision loss, and the second term causes the quantization oscillations in QAT.
-In order to emulate the effects of QAT, we propose a regularization term so that the training objective becomes:
-$\mathcal{L}(q(\mathbf{w}))=\mathcal{L}(\mathbf{w})+\mathcal{R_{\lambda}}(\mathbf{w})$ (22)
+In order to emulate the effects described in Section 3, we propose a regularization term so that the training objective becomes
+$\mathcal{L}(\mathbf{w})+\mathcal{R_{\lambda}}(\mathbf{w}),$ (18)
 where we let the regularization term be similar to the quadratic term in Eq. (9):
@@
 Here $\lambda\geq 0$ is a hyperparameter that controls the amount of regularization, $\ell$ ranges over the layers in the model and $i$ over the weights in each layer.
-Using the STE, ^$\frac{\hat{\partial}q}{\partial\mathbf{w}}=1$, we have the following expression for the gradient:
-^^$\frac{\hat{\partial}}{\hat{\partial}w^{\ell}_{i}}\mathcal{R_{\lambda}}(\mathbf{w})=\frac{\lambda}{n_{\ell}}\left(q(w^{\ell}_{i})-w^{\ell}_{i}\right)=-\frac{\lambda}{n_{\ell}}\mathbf{\Delta}(w_{i}^{\ell}).$ (24)
+In this term, we replaced the factor $x^{2}$ by a hyperparameter $\lambda$, since the precise expression of $x^{2}$ is specific to the model studied in Sec.
+3.
+We empirically find that this regularizer is sufficient to induce oscillations.
+The exploration of the design space of oscillation-inducing regularizers, including layer-dependent and/or adaptive scale factors, is left to future work.
+Using the STE, $\frac{\hat{\partial}q}{\partial\mathbf{w}}=1$, we have the following expression for the gradient:
+$\frac{\hat{\partial}}{\hat{\partial}w^{\ell}_{i}}\mathcal{R_{\lambda}}(\mathbf{w})=\frac{\lambda}{n_{\ell}}\left(q(w^{\ell}_{i})-w^{\ell}_{i}\right)=-\frac{\lambda}{n_{\ell}}\mathbf{\Delta}(w_{i}^{\ell}).$ (20)
 By the same reasoning as in Sec.
-4 this pulls the weight $w_{i}^{\ell}$ towards the quantization threshold and causes the gradient to “flip" as $w_{i}^{\ell}$ crosses the threshold.
+3 this pulls the weight $w_{i}^{\ell}$ towards the quantization threshold and causes the gradient to “flip” as $w_{i}^{\ell}$ crosses the threshold.
 We expect this to lead to oscillations based on the same mechanism as in the model from Sec.
-4.
+3.
 Figures 2 and  3 show the results of an experiment where we observe the weight distributions, and measured the oscillations, during training of a neural network (ResNet-18) with varying degrees of regularization, respectively.
-For comparison purposes the figures also shows the weight distributions and oscillations observed during training with QAT.
-Using the definition of an oscillation established in Nagel et al.
-(2022), we count an oscillation at epoch $i>1$ if $q(w_{t})\neq q(w_{t-1})$ and the direction of the change in the quantized space is opposite to that of the previous change.
-We note though that this method of counting misses the first threshold crossing during an oscillation A.5
-Our first observation is that QAT displays more oscillations – also seen as clustering around the quantization threshold in Fig.
-2-a) – than a baseline model without QAT or regularization (corresponding to $\lambda=0$ in Fig.
-3-b)) .
+For comparison, the figures also show the weight distributions and oscillations observed during training with QAT.
+Comparison t-score p-value
+$\lambda_{0}$ vs. $\lambda_{1}$ $25.13$ <0.001<0.001
+$\lambda_{0}$ vs. $\lambda_{10}$ $38.86$ <0.001<0.001
+$\lambda_{0}$ vs. QAT $18.83$ <0.001<0.001
+Table 1: Welch’s t-test comparing the per-weight oscillation count after 50 epochs of training, for the different cases in Fig.
+3.
+We note how the baseline $\lambda=0$ differs significantly from QAT and regularization induced oscillations with $\lambda=1,10$.
+Our first observation is that QAT displays more oscillations Fig.
+3-(a) than a baseline model without QAT or regularization (corresponding to $\lambda=0$ in Fig.
+3-(b)).
 As we increase $\lambda$ we observe that the number of oscillations as well as the clustering increases.
-This confirms that our regularizer can indeed induce oscillations similar to QAT during the training of deep neural networks.
+This confirms that the regularizer from Eq.
+(19) can indeed induce oscillations similar to QAT during the training of deep neural networks.
 At $\lambda=1$ (Fig.
-3-c)) the number of oscillations observed with our regularizer is similar to the behaviour of QAT, lending support to our hypothesis that the mechanism in (15) is indeed at the root of the oscillations observed when training neural networks with QAT.
+3-(c)) the number of oscillations observed with regularization is similar to the behavior of QAT, lending support to our hypothesis that the mechanism in Eq.
+(13) is indeed at the root of the oscillations observed when training neural networks with QAT.
+We use Welch’s $t$-test to test if the oscillation counts in Fig.
+3, which is the per-weight oscillation count after 50 epochs, are significantly different to the $\lambda=0$ baseline, reported in Table 1.
+Each of the pairwise comparison shows that the distribution of oscillations are significantly different.
 Figure 2: Weight distribution analysis of ResNet-18’s first convolutional layer after 50 epochs of training from scratch.
@@
 As $\lambda$ increases, we observe progressively stronger clustering of weights around quantization thresholds, illustrating the relationship between regularization strength and weight clustering.
-Figure 3: The plots show the distribution of weights with oscillation counts $>0$ when training with a) QAT and b)-d) our regularizer for different values of $\lambda$.
-Here $\lambda=0$ corresponds to a full precision model where our regularizer has no influence on training.
-The y-axis represents the percentage of total weights in the first convolutional layer of a ResNet-18 trained from scratch for 50 epochs, while the x-axis shows the oscillation count.
-Following the oscillation definition from Nagel et al.
-(2022), we count oscillations at each epoch during training.
+Figure 3: Distribution of weight oscillations.
+The plots show the distribution of weights with oscillation counts >0>0 when training with a) QAT and b)-d) the regularizer for different values of $\lambda$.
+Here $\lambda=0$ corresponds to a full-precision model where the regularizer has no influence on training.
+The y-axis represents the percentage of total weights in the first convolutional layer of a ResNet-18 trained from scratch for 50 epochs, while the x-axis shows the oscillation count after 50 epochs.
+Following the oscillation definition from nagel2022overcoming, we count oscillations at each epoch during training.
 The results demonstrate that QAT produces a significantly higher proportion of oscillating weights compared to $\lambda=0$.
 Furthermore, we observe that as we increase $\lambda$ a greater percentage of weights oscillates.
-Section 6 Experiments & Results
-In this section we empirically try to answer the question: is it sufficient to induce weight oscillations during training in order to get the benefits of QAT?
-We answer this question mostly affirmatively for ResNet and Vision Transformer architectures, based on the results of training ResNet-18 and Tiny ViT on the CIFAR-10 dataset.
-This is both in a training-from-scratch setting and when fine-tuning pretrained models.
+5 Experiments and Results
+In this section, we address the question: Is inducing weight oscillations during training sufficient to obtain the benefits of QAT?
+We answer this question affirmatively using empirical evidence based on the results of training ResNet-18 and Tiny Vit on the CIFAR-10 and Tiny-ImageNet datasets.
+This is both in the training-from-scratch (FS) setting and when fine-tuning (FT) pretrained models.
+In the following subsections we first describe the experimental setup, present accuracy results in FS and FT settings for models trained with different quantization levels for the quantizer in $\mathcal{R}_{\lambda}$ or QAT and finally, and report the cross-bit accuracy of the fine-tuned models.
+We train models at ternary (three possible values: -1, 0, 1), 3-bit and 4-bit.
+This is in line with contemporary research, where the emphasis lies on quantization at 4-bit and below, since the challenges of maintaining accuracy are more significant compared to quantization at higher bit widths.
+5.1 Experimental Setup
+We conducted our experiments using the CIFAR-10 krizhevsky2009learning and Tiny-ImageNet Le2015TinyIV datasets.
+We evaluated three architectures; a multi-layer perceptron with five hidden layers and 256 neurons per layer (MLP5), ResNet-18 he2016deep and Tiny ViT wu2022tinyvit.
+For each architecture we used the Adam optimizer kingma2014adam and tested multiple configurations: a baseline model to establish optimal floating-point (FP) accuracy and PTQ performance, a model with QAT, and a model with regularization using Eq.
+(19).
+The two latter configurations are trained using 3-bit and 4-bit quantizers.
 In all our experiments we use the regularizer $\mathcal{R}_{\lambda}$ defined in Eq.
-(23) to induce oscillations.
-In the following subsections we first describe the experimental setup, then we present the accuracy results from training-from-scratch and fine-tuning models trained with different quantization levels for the quantizer in $\mathcal{R}_{\lambda}$ or QAT and finally, we present the cross-bit accuracy of the fine-tuned models.
-We train models at ternary (3 possible values: -1, 0, 1), 3-bit and 4-bit.
-This is in line with contemporary research, where the emphasis lies on quantization at 4-bit and below since the challenges of maintaining accuracy are more significant compared to quantization at higher bit widths.
-Section 6.1 Experimental setup
-We conducted our experiments using the CIFAR-10 dataset Krizhevsky et al.
-(2009) without data augmentation.
-We evaluated three architectures; A multi-layer perceptron with 5 hidden layers and 256 neurons per layer (MLP5), ResNet-18 He et al.
-(2016) and Tiny Vision transformer (Tiny ViT) Wu et al.
-(2022).
-For each architecture we used the Adam optimizer Kingma (2014) and tested multiple configurations: A baseline model to establish optimal floating-point accuracy and post-training quantization (PTQ) performance, a model with QAT and a model with our approach.
-The two latter configurations are trained using a ternary, 3-bit, and 4-bit quantizer.
-Training from Scratch For the MLP5 architecture, we used a learning rate of $10^{-3}$ and regularization parameter $\lambda$=1.
+(19) to induce oscillations (marked as “Oscillations” in the results table).
+Training-From-Scratch (FS): For the MLP5 architecture, we used a learning rate of $10^{-3}$ and regularization parameter $\lambda$=1.
 The ResNet-18 was trained with a learning rate of $10^{-3}$ and $\lambda$=0.75 (see Appx.
@@
 For quantized models, we monitored the quantized validation accuracy at the target bit precision, while for the baseline, we tracked floating-point accuracy.
-Fine-tuning We fine-tuned two ImageNet-1k Deng et al.
-(2009) pre-trained models on CIFAR-10: a Tiny ViT (learning rate: $10^{-4}$, $\lambda$=1) and a ResNet-18 (learning rate: $10^{-3}$, $\lambda$=1).
-To maintain compatibility with the pre-trained architectures, we upsampled CIFAR-10 images to $224\times 224$ pixels.
-The $\lambda$ parameter selection process for Tiny ViT is detailed in Appx.
-A.2.
-Fine-tuning continued for up to 200 epochs, with early stopping after 30 epochs without improvement, using the same accuracy metrics as training from scratch.
-Quantization We implemented weight quantization using a per-tensor uniform symmetric quantizer as defined in Eq.
-1.
+Fine-tuning (FT): We fine-tuned two ImageNet-1k deng2009imagenet pre-trained models on CIFAR-10 and Tiny-ImageNet: a Tiny ViT (learning rate: $10^{-4}$, $\lambda\in\{1,0.75,0.5\}$ depending on the bit) and a ResNet-18 (learning rate: $10^{-3}$, $\lambda\in\{1,0.75,0.5\}$ depending on the bit).
+To maintain compatibility with the pre-trained architectures, we up-sampled both CIFAR-10 and Tiny-Imagenet images to $224\times 224$ pixels.
+The $\lambda$ parameter selection process for Tiny ViT is detailed in Appendix A.2.
+Fine-tuning continued for up to 200 epochs on CIFAR-10 and 50 epochs for Tiny-ImageNet, with early stopping after 30 epochs without improvement, using the same accuracy metrics as training from scratch.
+Quantization: We implemented weight quantization using a per-tensor uniform symmetric quantizer as defined in Eq.
+(1).
+PTQ is applied in its most minimal form, by simply quantizing the weights without any calibration.
+QAT is used as defined in Eq.(5).
 The quantization range was determined by computing minimum and maximum values per layer.
@@
 And lastly for the MLP5 model all layers were quantized.
-Section 6.2 Training-from-scratch
-Table 1 shows the results from training an MLP and ResNet-18 from scratch on the CIFAR-10 dataset.
-Our regularization method (OsciQuant) demonstrates improvements compared to the PTQ baseline from ternary quantization.
+For Tiny-ImageNet models are trained at 3 and 4-bit precision only.
+5.2 Results
+The performance on the two datasets in training-from-scratch and fine-tuning settings is presented in the following sections, along with the observations about cross-bit generalization.
+5.2.1 Performance on CIFAR-10
+Training from Scratch:
+Table 2 (A) shows the results from training an MLP and ResNet-18 from scratch on the CIFAR-10 dataset.
+Doing only regularization with Eq.
+(19) demonstrates improvements compared to the PTQ baseline.
 More importantly, it also matches the performance of QAT at bit widths of 3 and 4.
-For both models we see that at 3-bit and 4-bit, our method exhibits similar performance as QAT but with less variability, while not differing significantly in the average number of training epochs required.
-With both models, QAT and OsciQuant are competitive with the full-precision baseline, although we observe an increased number of training epochs.
-Notably, both OsciQuant and QAT significantly outperform PTQ when applied to the full precision baseline.
-Model Quantization method Accuracy Mean Epochs
-MLP5 Baseline FP32 51.43 $\pm$ 0.39 14
-Ternary PTQ 10.00 $\pm$ 0.02 14
-Ternary QAT 49.20 $\pm$ 1.34 24
-Ternary OsciQuant 36.49 $\pm$ 0.51 14
-3-bit PTQ 20.97 $\pm$ 5.64 14
-3-bit QAT 50.53 $\pm$ 1.43 33
-3-bit OsciQuant 48.48 $\pm$ 0.29 15
-4-bit PTQ 46.50 $\pm$ 0.76 14
-4-bit QAT 51.39 $\pm$ 0.60 26
-4-bit OsciQuant 50.72 $\pm$ 0.47 19
-ResNet-18 Baseline FP32 83.26 $\pm$ 1.07 24
-Ternary PTQ 10.00 $\pm$ 0.01 24
-Ternary QAT 79.62 $\pm$ 6.42 42
-Ternary OsciQuant 61.5 $\pm$ 1.82 56
-3-bit PTQ 77.79 $\pm$ 4.0 24
-3-bit QAT 82.51 $\pm$ 2.14 37
-3-bit OsciQuant 81.77 $\pm$ 0.46 41
-4-bit PTQ 82.11 $\pm$ 1.21 24
-4-bit QAT 82.66 $\pm$ 2.57 28
-4-bit OsciQuant 83.74 $\pm$ 0.59 32
-Table 1: Comparison of accuracy when training from scratch on CIFAR-10.
-Results show classification accuracy and mean training epochs for MLP5 and ResNet-18 across different quantization approaches and bit-widths.
-Results is means and standard deviations over 5 random seeds.
-Model Quantization method Accuracy Mean Epochs
-ResNet-18 Baseline FP32 88.50 $\pm$ 0.64 4
-Ternary PTQ 10.01 $\pm$ 0.01 4
-Ternary QAT 77.02 $\pm$ 7.57 47
-Ternary OsciQuant 44.59 $\pm$ 3.30 35
-3-bit PTQ 10.28 $\pm$ 0.48 4
-3-bit QAT 85.69 $\pm$ 1.83 25
-3-bit OsciQuant 84.94 $\pm$ 1.59 27
-4-bit PTQ 35.56 $\pm$ 9.05 4
-4-bit QAT 87.71 $\pm$ 1.14 26
-4-bit OsciQuant 87.08 $\pm$ 0.72 24
-Tiny ViT Baseline FP32 96.11 $\pm$ 0.31 6
-Ternary PTQ 9.39 $\pm$ 1.11 6
-Ternary QAT 73.53 $\pm$ 0.77 140
-Ternary OsciQuant 13.51 $\pm$ 1.32 28
-3-bit PTQ 11.56 $\pm$ 1.99 6
-3-bit QAT 88.13 $\pm$ 0.60 131
-3-bit OsciQuant 88.68 $\pm$ 1.08 108
-4-bit PTQ 21.57 $\pm$ 5.33 6
-4-bit QAT 94.96 $\pm$ 0.33 57
-4-bit OsciQuant 94.82 $\pm$ 0.51 90
-Table 2: Comparison of accuracy when fine-tuning on models pre-trained on ImageNet-1k.
-Results show classification accuracy and mean training epochs for MLP5 and ResNet-18 across different quantization approaches and bit-widths.
-Results is means and standard deviations over 5 random seeds.
-Section 6.3 Fine-tuning
-Table 2 summarizes the test accuracies for fine-tuning using our OsciQuant method and QAT on ResNet-18 and Tiny ViT architectures.
-The observations are roughly in line with the results observed for training from scratch in the previous section with the exception of the number of epochs required for fine-tuning.
-On the ResNet architecture both QAT and our model train for significantly longer than the full precision baseline.
-As is the case for training from scratch, we see an increase in ternary performance compared to PTQ, but QAT still ourperforms our method in the ternary setting.
-Our regularization and QAT show comparable performance when quantized at 3 bits and 4 bits, while achieving test accuracy close to the full precision model at 4-bits.
-The general trend regarding accuracy is identical for the vision transformer experiments, while we again note the high number of epochs require for both methods when fine-tuning, compared to the full precision baseline.
-Section 6.4 Robustness to cross-bit quantization
-As described above, the goal of our proposed regularization term is to train a model that maintains performance after quantization.
-Since the regularization term involves a quantization operator, we need to choose the quantization level in the regularization term.
-In this experiment we evaluated the robustness of our method and QAT towards quantization at levels different from the ones used during training.
-For OsciQuant, we applied a regularization term with the training bit width during training and applied PTQ after training finished at a different quantization level.
+For both models we see that at 3-bit and 4-bit, using the $\mathcal{R_{\lambda}}$ regularizer from Eq.
+(19) exhibits similar performance as QAT but with less variability.
+With both models, QAT and $\mathcal{R_{\lambda}}$ regularization are competitive with the full-precision baseline.
+Notably, both $\mathcal{R_{\lambda}}$ regularization and QAT significantly outperform PTQ when applied to the full-precision baseline.
+Quantization MLP5 (FS) ResNet-18 (FS) Tiny ViT (FT) ResNet-18 (FT)
+Baseline FP32 51.43 $\pm$ 0.39 83.26 $\pm$ 1.07 96.11 $\pm$ 0.31 88.50 $\pm$ 0.64
+3-bit PTQ 20.97 $\pm$ 5.64 77.79 $\pm$ 4.00 11.56 $\pm$ 1.99 10.28 $\pm$ 0.48
+3-bit QAT 50.53 $\pm$ 1.43 82.51 $\pm$ 2.14 88.13 $\pm$ 0.60 85.69 $\pm$ 1.83
+3-bit Oscillations 48.48 $\pm$ 0.29 81.77 $\pm$ 0.46 88.68 $\pm$ 1.08 84.94 $\pm$ 1.59
+4-bit PTQ 46.50 $\pm$ 0.76 82.11 $\pm$ 1.21 21.57 $\pm$ 5.33 35.56 $\pm$ 9.05
+4-bit QAT 51.39 $\pm$ 0.60 82.66 $\pm$ 2.57 94.96 $\pm$ 0.33 87.71 $\pm$ 1.14
+4-bit Oscillations 50.72 $\pm$ 0.47 83.74 $\pm$ 0.59 94.82 $\pm$ 0.51 87.08 $\pm$ 0.72
+Table 2: Performance comparison on CIFAR-10 dataset.
+Results show classification accuracy for MLP5, ResNet-18, and Tiny ViT across different quantization approaches and bit-widths.
+Models trained from scratch are marked FS, and fine-tuning experiments are marked FT.
+FT experiments are based on models pre-trained on ImageNet-1k.
+In all cases oscillations followed by quantization of the weights matches QAT accuracy.
+Results are means and standard deviations over 5 random seeds.
+PTQ results are from the FP32 baseline.
+Fine-tuning:
+Table 2 (B) summarizes the test accuracies for fine-tuning using our $\mathcal{R_{\lambda}}$ regularization and QAT on ResNet-18 and Tiny ViT architectures with CIFAR-10 and Tiny-ImageNet.
+The observations are roughly in line with the results observed for training from scratch in the previous section.
+For CIFAR-10 as with the case for training from scratch, with both ResNet-18 and Tiny Vit, we see an increase in performance compared to PTQ when using the regularization in Eq.
+(19).
+Regularization with $\mathcal{R_{\lambda}}$ and QAT show comparable performance when quantized at 3 bits and 4 bits, while achieving test accuracy close to the full-precision model at 4-bits.
+Performance comparison and related discussions for ternary quantization are presented in Appendix A.3.
+Robustness to Cross-bit Quantization:
+Model Train bit $\downarrow$ / Eval. bit $\rightarrow$ FP32 Ternary 3-bit 4-bit 8-bit
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 3: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on CIFAR-10.
+Grey background is the target-bit accuracy.
+Models are trained using different quantization methods (QAT and ours) and bit-widths (3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32.
+Results are means and standard deviations over 5 random seeds.
+All significant differences between QAT and Oscillations are shown in bold face.
+In this experiment we evaluated the robustness of oscillations-only and QAT towards quantization at bit widths different from the ones used during training.
+When using the $\mathcal{R_{\lambda}}$ regularization approach, we applied the regularization term with the training bit width during training and applied PTQ after training at a different quantization level.
 For QAT we trained using the training bit width and afterwards applied PTQ to the latent weights.
 For each method we also evaluated the corresponding model without PTQ, directly using the latent weights for inference (reported as FP32).
-Table 3 shows the results from the experiment.
-A first observation is that the models produced by our method consistently achieve nearly full-precision accuracy when quantized at 8-bit or when used without quantization, irrespective of the quantization level used during training.
+In Table 3 the results from the experiment are reported.
+A first observation is that the models produced by $\mathcal{R_{\lambda}}$ regularization consistently achieve nearly full-precision accuracy when quantized at 8-bit or when used without quantization, irrespective of the quantization level used during training.
 This contrasts with QAT, which produces a viable 8-bit or full-precision model only when trained with at least 4-bit.
-Furthermore we see that our method mostly maintains performance when trained at 3 or 4-bit and quantized at bit level of 3 or 4-bit.
+Furthermore we see that regularizing using Eq.
+(19) mostly maintains performance when trained at 3 or 4-bit and quantized at bit level of 3 or 4-bit.
 QAT also achieves this for Tiny ViT but for ResNet, the accuracy of QAT trained at 3-bit and quantized at other bit widths is barely above random guessing.
-Regarding training with ternary quantization, we see that our method produces models that achieve near full precision performance for ResNet when quantized at 3-bit or higher.
+Regarding training with ternary quantization, we see that $\mathcal{R_{\lambda}}$ regularization produces models that achieve near full-precision performance for ResNet when quantized at 3-bit or higher.
 Ternary training for ViT is somewhat peculiar in that it fails to produce a model that is viable when quantized to ternary, whereas the performance of the resulting models starts to show a high level of variability at 4-bit and finally reaches close to full-precision accuracy at 8-bit.
 In contrast, for both ResNet and ViT, the performance of QAT degrades completely to random guessing when trained with ternary quantization and evaluated at any other quantization level.
-Model Train bit $\downarrow$ / Eval. bit $\rightarrow$ FP32 Ternary 3-bit 4-bit 8-bit
-ResNet-18 Baseline (PTQ) 88.50 $\pm$ 0.64 10.01 $\pm$ 0.01 10.28 $\pm$ 0.48 35.56 $\pm$ 9.05 88.45 $\pm$ 0.64
-Ternary QAT 10.39 $\pm$ 0.71 77.02 $\pm$ 7.57 9.75 $\pm$ 0.77 10.03 $\pm$ 0.51 10.35 $\pm$ 0.63
-Ternary OsciQuant 87.44 $\pm$ 0.56 44.59 $\pm$ 3.30 85.42 $\pm$ 1.13 87.03 $\pm$ 0.65 87.42 $\pm$ 0.56
-3-bit QAT 16.89 $\pm$ 4.97 10.01 $\pm$ 0.04 85.69 $\pm$ 1.83 17.42 $\pm$ 4.96 16.56 $\pm$ 4.32
-3-bit OsciQuant 87.86 $\pm$ 0.42 20.19 $\pm$ 10.74 84.94 $\pm$ 1.59 87.56 $\pm$ 0.38 87.86 $\pm$ 0.42
-4-bit QAT 87.75 $\pm$ 1.13 10.13 $\pm$ 0.29 82.08 $\pm$ 6.25 87.71 $\pm$ 1.14 87.76 $\pm$ 1.12
-4-bit OsciQuant 87.85 $\pm$ 0.49 11.91 $\pm$ 0.87 85.57 $\pm$ 1.10 87.08 $\pm$ 0.72 87.87 $\pm$ 0.49
-Tiny ViT Baseline (PTQ) 96.11 $\pm$ 0.31 9.39 $\pm$ 1.11 11.56 $\pm$ 1.99 21.57 $\pm$ 5.33 96.03 $\pm$ 0.34
-Ternary QAT 10.62 $\pm$ 1.29 73.53 $\pm$ 0.77 11.52 $\pm$ 1.82 11.13 $\pm$ 1.75 10.61 $\pm$ 1.26
-Ternary OsciQuant 95.79 $\pm$ 0.58 13.51 $\pm$ 1.32 12.53 $\pm$ 3.66 54.93 $\pm$ 27.32 95.76 $\pm$ 0.59
-3-bit QAT 86.94 $\pm$ 0.91 19.78 $\pm$ 6.04 88.13 $\pm$ 0.60 86.69 $\pm$ 0.62 86.95 $\pm$ 0.89
-3-bit OsciQuant 96.47 $\pm$ 0.11 9.48 $\pm$ 1.64 88.68 $\pm$ 1.08 95.35 $\pm$ 0.18 96.50 $\pm$ 0.11
-4-bit QAT 95.14 $\pm$ 0.29 $1$$\pm$ 1.84 59.86 $\pm$ 19.95 94.96 $\pm$ 0.33 95.13 $\pm$ 0.28
-4-bit OsciQuant 96.54 $\pm$ 0.09 11.90 $\pm$ 1.29 70.23 $\pm$ 12.75 94.82 $\pm$ 0.51 96.55 $\pm$ 0.09
-Table 3: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on CIFAR-10.
+5.2.2 Performance on Tiny-ImageNet
+Fine-tuning:
+Table 4 summarizes the test accuracies for the Tiny-ImageNet dataset.
+Here we observe the same tendency as with CIFAR-10; oscillations provides a significant increase in accuracy compared to the PTQ baseline.
+While for the Tiny ViT model $\mathcal{R_{\lambda}}$ regularization is sufficient to recover the quantized accuracy of QAT in both the 3 and 4-bit case, for ResNet18 there is a degradation in accuracy at 3-bit.
+Quantization Tiny ViT (FT) ResNet-18 (FT)
+Baseline FP32 67.17 $\pm$ 0.67 62.93 $\pm$ 0.55
+3-bit PTQ 0.58 $\pm$ 0.16 0.51 $\pm$ 0.03
+3-bit QAT 44.29 $\pm$ 0.49 54.08 $\pm$ 0.52
+3-bit Oscillations 44.62 $\pm$ 2.47 49.34 $\pm$ 0.76
+4-bit PTQ 11.02 $\pm$ 2.11 20.02 $\pm$ 4.80
+4-bit QAT 60.61 $\pm$ 0.16 58.31 $\pm$ 0.19
+4-bit Oscillations 60.54 $\pm$ 0.37 57.26 $\pm$ 0.33
+Table 4: Accuracy on Tiny-ImageNet dataset.
+Mean and standard deviation is over 3 runs.
+The models is fine-tuned for 50 epochs on the pretrained ImageNet models.
+PTQ results is from the FP32 baseline.
+In both the 3 and 4 bit case, oscillations followed by quantization of the weights matches QAT and is noticeably above the PTQ baseline which has neither oscillations nor QAT.
+Robustness to Cross-bit Quantization:
+In Table  5 we see the cross-bit results from the Tiny-ImageNet experiments.
+As with CIFAR-10 we note that the models produced by $\mathcal{R_{\lambda}}$ regularization achieves a better cross-bit performance at bits higher than the target bit.
+Though we do note a change in the cross-bit behavior; the cross-bit results for 3 and 4-bit is generally lower and not as close tot he FP baseline as in the CIFAR-10 case, yet still there is a significant difference between QAT and $\mathcal{R_{\lambda}}$ regularization.
+Model Method $\downarrow$ / Eval. bit $\rightarrow$ FP32 Ternary 3-bit 4-bit 8-bit
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 5: Cross-bit evaluation of pre-trained ImageNet-1k models fine-tuned on Tiny-ImageNet.
 Grey background is the target-bit accuracy.
-Models are trained using different quantization methods (QAT and ours) and bit-widths (ternary, 3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32.
-The grey diagonal shows the results for the bit used during training.
+Models are trained using different quantization methods (QAT and ours) and bit-widths (3-bit, and 4-bit), then evaluated across various bit-widths ranging from ternary to FP32.
 Results are means and standard deviations over 5 random seeds.
-All significant differences between QAT and OsciQuant are shown in bold face.
-Section 7 Discussion
-We have shown that training with weight oscillations induced via regularization is sufficient in most cases to maintain performance after quantization for ResNet and Tiny ViT.
+All significant differences between QAT and Oscillations are shown in bold face.
+6 Discussion
+We have shown that training with weight oscillations induced via $\mathcal{R_{\lambda}}$ regularization is sufficient in most cases to maintain performance after quantization for ResNet and Tiny ViT.
+The primary effect of our regularizer is to push weights toward quantization thresholds, which results in weights that are robust to quantization.
+This “threshold-pushing” dynamic naturally leads to two observable phenomena: weight clustering and oscillations.
 This begs the question whether weight oscillations are also a necessary part of the QAT training process.
 Indeed, some previous work already points towards this.
-There are examples claiming that both dampening and/or freezing of oscillations too early in the training process is detrimental to performance after quantization Nagel et al.
-(2022); Han et al.
-(2021).
-And in other case presented in Liu et al.
-(2023), freezing only the low frequency oscillating weights improves performance.
+There are examples claiming that both dampening and/or freezing of oscillations too early in the training process is detrimental to performance after quantization nagel2022overcoming; ImprovingLowBit.
+And in other case presented in vitoscillations, freezing only the low frequency oscillating weights improves performance.
 This suggests that weight oscillations are both a necessary and sufficient part of QAT, at least in the early phases of the training process.
-This further supports our hypothesis that oscillations in QAT have a positive effect on quantization robustness.
-Additionally, there might be further benefits to our regularization approach compared to QAT.
-Our method aims to isolate this crucial part of the training process.
-This is arguably a more principled approach compared to QAT, where quantization during training combined with STE can lead to a number of side-effects beyond oscillations, which can be highly non-intuitive.
-We present a simple example in the Appendix Sec.
-A.1 where replacing a single scalar weight by a product of two scalar weights leads to a non-trivial change in training dynamics when using QAT with the STE.
-On the other hand, while it is not clear what the additional effects are during QAT, we do note two consistent deviations from the QAT performance when using our regularization method: QAT outperforms regularization at ternary quantization, whereas our regularization method outperforms QAT in cross-bit accuracy for the ternary and 3-bit case.
-In A.4, we see how it seems that the cross-bit performance for QAT is upper-bounded by the target-bit performance, which might explain the subpar QAT performance at cross-bit compared to our regularization method which seems bounded by the full precision accuracy.
-Additionally we can note that while it is stated in Alizadeh et al.
-(2020); Chmiel et al.
-(2020) that QAT is not robust to cross-bit quantization, A.4 shows that for some cases the robustness is tied closely to how long the model is trained after the target bit accuracy has converged.
-Finally we note in A.2 that in the ResNet-18 model, we see similar results for the hyperparameter sweep for different $\lambda$s, which might suggest that the key for robustness is the presence of oscillations and not their precise nature.
-Limitations In our experiments we observed that the robustness to cross-bit quantization improves in later training epochs.
-In order to further improve robustness one might consider an early stopping criterion that evaluates the performance on cross-bit quantization, which was not done in this work.
-The same approach could also increase cross-bit quantization robustness of QAT although to a lesser degree than for our method.
-We performed our experiments on the CIFAR-10 dataset which might make it more difficult to compare our results with other published works that provide benchmark results for other datasets such as ImageNet-1k.
-Section 8 Conclusion
-Based on the analysis of a toy model we proposed the hypothesis that weight oscillations during training in deep neural networks make the model robust to quantization.
-In Sections 4 and 5 we explain on a toy model how training with QAT and STE leads to oscillations and propose a regularizer that encourages this oscillating behaviour.
-We confirm that as we increase the strength of the regularization, we empirically observe the appearance of clustering together with oscillations.
-Finally we experimentally confirm that the regularizer indeed leads to consistent robustness towards quantization for quantization levels above ternary.
-Our regularization method achieves comparable performance to QAT above ternary quantization when quantizing to the target-bit seen during optimizing and shows increased robustness compared to QAT in cross-bit quantization with bits greater than the target-bit used in the quantizer during training.
-All this being evidence of our hypothesis.
-Our insights on weight oscillations and their role in quantization robustness open new horizons for model quantization approaches.
-Our regularization method especially creates interesting possibilities for cross-bit robustness, potentially making our regularization method more appealing than QAT when the goal is to deploy or relase a single set of weights that works across different bit widths or maybe even quantizers.
+This further supports our hypothesis that oscillations in QAT have a positive effect on quantization robustness overall.
+Yet we do note deviations from QAT when regularizing with Eq.
+(19): QAT outperforms $\mathcal{R_{\lambda}}$ regularization at ternary quantization (Appendix A.3), whereas $\mathcal{R_{\lambda}}$ regularization outperforms QAT in cross-bit accuracy for the ternary and 3-bit case.
+In A.5, we see how it seems that the cross-bit performance for QAT is upper-bounded by the target-bit performance, which might explain the subpar QAT performance at cross-bit compared to $\mathcal{R_{\lambda}}$ regularization which seems bounded by the full-precision accuracy.
+Limitations and Future Work
+Our theoretical analysis was performed using the toy model in Section 3, and the regularization term is motivated using this analysis.
+We expect other effects that are not entirely captured by this analysis to be part of QAT.
+This is explored further in Appendix A.1, where we show how the second term in Eq.
+(9) is not zero in the gradient when there are multiple layers.
+The second term in Eq.
+(9) is closely related to the oscillations-dampening methods such as the one presented in Equation 6 in nagel2022overcoming, which works by adding a linear pull towards the quantization level.
+In Appendix A.1 we show how for the linear case with two weights, the second term is no longer zero in the gradient and as such QAT then consists of two components; one that creates oscillations and one that dampens them.
+From this we can see that oscillations alone do not capture the full dynamics of QAT as analyzed in linear models.
+Broader Impact:
+The work presented here investigates the oscillations that arise when preparing deep neural networks for quantization using QAT.
+By characterizing these oscillations, we gain deeper insight into the dynamics of QAT, which can be leveraged to enhance its effectiveness and, consequently, the efficiency of quantized models.
+Improved efficiency translates into reductions in memory, computation, and energy demands, facilitating the deployment of high-performing quantized networks on low-power edge devices.
+Some of these applications can have broader positive impact (for example, medical electronics) and also negative impact (for example, surveillance systems or weapons).
+7 Conclusions
+Based on the analysis of a linear model we hypothesized that weight oscillations during training in deep neural networks make the model robust to quantization akin to QAT.
+In Sections 3 and 4 we explain how training with QAT and STE leads to oscillations and propose a regularizer that encourages this oscillating behavior.
+We confirm that as we increase the strength of the regularization, and empirically observe the appearance of clustering together with oscillations.
+Finally, we experimentally confirm that the regularizer indeed leads to consistent robustness towards quantization for 3-bit and 4-bit quantization levels.
+Our oscillations by regularization approach achieves comparable performance to QAT above ternary quantization when quantizing to the target-bit seen during the optimization.
+Furthermore, we also observe that it shows increased robustness compared to QAT in cross-bit quantization with quantization levels higher than the target-bit used in the quantizer during training.
+All this is evidence in favor of our hypothesis.
+Our insights on weight oscillations and their role in quantization robustness open new horizons for model quantization approaches which usually build on the idea of aligning weights at quantization levels – the opposite of what seems to be the core dynamic in QAT.
+The regularization approach especially creates interesting possibilities for cross-bit robustness, potentially making the regularization method more appealing than QAT when the goal is to deploy or release a single set of model weights that can work across different bit widths or maybe even quantizers.
 While the regularizer used in our experiments should be viewed as an initial step, we expect that quantization robustness could be further improved by developing oscillation-inducing methods that are adaptive to different learning rates, layer statistics or phases of the training process.
-Acknowledgments
-Authors thank Tong Chen, Jákup Svøðstein and Sebastian Hammer Eliassen for useful discussions.
+Acknowledgments:
+Authors thank Tong Chen, Jákup Svøðstein, Sophia Wilson, Sebastian Hammer Eliassen, Shivam Adarsh and members of SAINTS Lab for useful discussions throughout.
 JW, BP and RS are partly funded by European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No.
@@
 Appendix A Appendix
-Section A.1 2-layer with single weights
-Figure 4: We repeat the toy model experiments, but this time with 2 weights, taking into account that the linear term is no longer 0 in the gradient.
-We notice at epoch 15 and 18 where the prediction of the quantized model is greater than y, the effect of the terms flip for $w_{2}$.
-Consider a linear model $f(x)=w_{2}w_{1}x$, with $w_{1},w_{2}$, input $x$, and target $y\in\mathbb{R}$.
+A.1 Two-layer with Single Weights
+Consider a linear model $f(x)=w_{2}w_{1}x$, with weights $w_{1},w_{2}$, input $x$, and target $y\in\mathbb{R}$.
 The quantized version of this model is defined as $f_{q}(x)=q(w_{2})q(w_{1})x$, where $q(\cdot)$ is the quantizer from Eq.
-1.
+Equation˜1.
 The quadratic loss for the model is given by
@@
 The loss difference decomposes into:
-$\underbrace{\frac{1}{2}x^{2}\Bigl{(}q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr{)}}_{\text{quadratic term}}\quad+\quad\underbrace{yx\Bigl{(}w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr{)}}_{\text{linear term}}$
-Taking the derivative of $\mathcal{L}$ with respect to $w_{1}$:
+$\underbrace{\frac{1}{2}x^{2}\Bigl(q(w_{2})^{2}q(w_{1})^{2}-w_{2}^{2}w_{1}^{2}\Bigr)}_{\text{Quadratic term (Oscillator)}}\quad+\quad\underbrace{yx\Bigl(w_{2}w_{1}-q(w_{2})q(w_{1})\Bigr)}_{\text{Linear term (Oscillation Dampener)}}$ (25)
+Taking the derivative of $\mathcal{L}$ w.r.t $w_{1}$:
 $\frac{\partial\delta_{\mathcal{L}}}{\partial w_{1}}$ $=\frac{\partial}{\partial w_{1}}\Bigl(\mathcal{L}(f_{q}(x))-\mathcal{L}(f(x))\Bigr)$ (26)
@@
 $\frac{\partial\hat{\delta}_{\mathcal{L}}}{\partial w_{1}}=x^{2}\Bigl[q(w_{2})^{2}q(w_{1})-w_{2}^{2}w_{1}\Bigr]+yx\Bigl[w_{2}-q(w_{2})\Bigr]$ (29)
-We note that the linear term is no longer zero in the gradient and thus for a model consisting of 2 single weight layers we see that there is additional effects from QAT other than oscillations.
-Additionally because of the non-linearity of the rounding operation, even with the absence of a non-linear activation function, we can no longer reduce the model to a single weight.
-Section A.2 Hyperparameters
+We note that the linear term is no longer zero in the gradient and thus for a model consisting of two single weight layers we see that there are additional effects from QAT other than oscillations.
+Furthermore, we would like to clarify that even though the two layered model under consideration does not have explicit non-linear activation functions, it cannot be reduced to a single layer model.
+This is because of the non-linearity of the rounding operation.
+Figure 4: We repeat the toy model experiments, but this time with two weights, taking into account that the linear term is no longer zero in the gradient.
+We notice at epoch 15 and 18 where the prediction of the quantized model is greater than y, the effect of the terms flip for $w_{2}$.
+A.2 Hyperparameters
 A.2.1 ResNet-18
+[…]
+Figure 5: Mean over 3 runs of the best test accuracy for different lambdas.
+Fine-tuning a pretrained ResNet-18 on CIFAR-10 for 50 epochs.
+Quantizer is set to 3-bit and $10^{-3}$ learning rate and 100% of the training data is used.
 $\lambda$ 3-bit (%) Ternary (%)
@@
 1.75 69.92 $\pm$ 0.11 47.02 $\pm$ 4.19
-Figure 5: Mean over 3 runs of the best validation accuracy for different lambdas.
-Training a ResNet-18 from scratch.
-Both ternary and 3-bit is at $10^{-3}$ learning rate and 50% of the data used for training.
-The plot shows three learning rates, where we for each learning ratue evaluate with the $\lambda$s in the rhs.
-table.
-The colored background covers the range between the maximum and minimum value of the quantized validation accuracy with the given $\lambda$s.
-In Fig.
-5 we see the results of a hyperparameter search over different learning rates and $\lambda$s for a ResNet-18 model.
-There is a clear trend of seeing the best performance at a learning rate of $10^{-3}$.
-We note that interestingly there is a comparable performance for a wide range of $\lambda$s, indicating that it is the presence of oscillations which is important for quantization robustness, and not the exact frequency of oscillations.
+[…]
 A.2.2 Tiny ViT
@@
 5.0 - 14.20
-Figure 6: Validation accuracy at different $\lambda$ values and the corresponding best validation accuracies for 3-bit and 2-bit configurations for a single run.
-Learning rate is set to 1e-4 for fine-tuning.
-For the 2-bit we test higher $\lambda$ but still see no improvemenet in accuracy.
-We note how all the $\lambda$s lies close to each other, except for the low of $10^{-2}$
-Fig. 6 We note how also the Tiny Vit seems to allow for a wide range of $\lambda$s even though we this time note that $\lambda=1$ performs significantly better than the others.
-Section A.3 Epochs and cross-bit robustness
+[…]
+[…]
+A.3 Ternary Quantization
+Performance comparison for the ternary quantization for different models on CIFAR-10 is reported in Table 6.
+While both QAT and oscillations improve the PTQ baseline significantly, oscillations degrade compared to QAT, especially for the Tiny ViT.
+This is in line with previous literature vitoscillations, where transformers have been identified as especially sensitive to oscillations.
+If oscillations are fundamental to QAT, this still leaves the question of why QAT achieves superior ternary performance.
+Appendix A.1 offers an indication of the underlying reason.
+Looking at the gradient of the two-layer toy model, in Eq.
+25 we note that the second term is no longer zero in the gradient.
+This term bears close resemblance to existing formulations of oscillation dampeners, such as in Eq.
+6 of nagel2022overcoming.
+These works by pushing weights towards their nearest quantization level, thereby dampening or nullifying the effect of any oscillator (which pushes weights towards their nearest threshold).
+We therefore speculate that the main component of QAT is still oscillations, but that QAT also has an inherent dampening mechanism.
+Given that in a uniform quantizer, the quantization error increases exponentially as we decrease the bits in the quantizer, ternary quantization’s large error magnitude makes the absence of dampening particularly detrimental, resulting in sub-optimal quantized performance compared to QAT.
+Quantization MLP5 (FS) ResNet-18 (FS) ResNet-18 (FT) Tiny ViT (FT)
+Baseline FP32 51.43 $\pm$ 0.39 83.26 $\pm$ 1.07 83.26 $\pm$ 1.07 96.11 $\pm$ 0.31
+Ternary PTQ 10.00 $\pm$ 0.02 10.00 $\pm$ 0.01 10.00 $\pm$ 0.01 9.39 $\pm$ 1.11
+Ternary QAT 49.20 $\pm$ 1.34 79.62 $\pm$ 6.42 77.02 $\pm$ 7.57 73.53 $\pm$ 0.77
+Ternary Oscillations 36.49 $\pm$ 0.51 61.50 $\pm$ 1.82 44.59 $\pm$ 3.30 13.51 $\pm$ 1.32
+Table 6: Performance comparison with ternary quantization on CIFAR-10 dataset.
+Mean and standard deviation is over 3 runs.
+The models is fine-tuned for 50 epochs on the pretrained ImageNet models.
+PTQ results is from the FP32 baseline.
+For both oscillations only and QAT we see a significant improvement over the PTQ baseline.
+Yet oscillations degrade significantly compared to QAT, especially for the Tiny Vit.
+FS: Train from scratch.
+FT: Fine-tuned.
+A.4 Epochs and cross-bit robustness
 Figure 8: Left is the validation accuracy during training of a ViT with QAT at different bits, right is for our regularization.
@@
 We note how QAT converges first for the target-bit and then over time also converges for the 4 and 8-bit.
-Additionally we see that QAT seems upper-bounded by the target-bit performance, while this is not the case for our metho.
-Fig.
-7,
-Section A.4 Convergence behaviour of Tiny ViT
-Figure 8: Regularization with a 3-bit quantizer on a Tiny ViT.
-We note the peculiar behaviour of the orange line, which is the validation accuracy on the target-bit performance.
+Additionally we see that QAT seems upper-bounded by the target-bit performance, while this is not the case for oscillations only as shown in Fig.
+8.
+A.5 Convergence behavior during oscillations-only optimization
+Figure 9: (right) Convergence behavior of ResNet-18.
+(left) Convergence behavior of a Tiny ViT with regularization with a 3-bit quantizer.
+Note the peculiar behavior of the orange line, which is the validation accuracy on the target-bit performance.
 The performances cycles between $\approx 90\%$ and $10\%$, while the full precision accuracy (the model evaluated without quantized weights) stays some-what stable.
 Fig.
-8 shows the convergence behaviour of the full precision weights and the quantized weights at target-bit.
-We note how the Tiny ViT displays a peculiar convergence behaviour, where the accuracy will break, only to go up again.
-In the Tiny Vit model we quantize the self-attention layers, it is already noted in  Liu et al.
-[2023] that ViTs are especially vulnerable to quantization of the query and key of a self-attention layer, which might be related to the convergence behaviour we see.
-Section A.5 Counting oscillations
-Figure 9: The plots shows the weights with a total oscillation count $>0$ in a baseline model, a QAT model and a model regularized with $\lambda=1$ respectively.
-The y-axis is the percentage of the total weights in the first convolutional layer in a ResNet-18 trained from scratch for 50 epochs.
-For the baseline model we log the full precision weights at each epoch and then apply the quantizer after training, where as for the regularized model we simply log both the full precision and quantized weights at each epoch.
-Using the counting method proposed by Nagel et al.
-[2022], we notice the biased odd distribution of the bins.
-In the plot used for the main paper, we count each bin the histogram in iterations of 2.
-Additionally we note that the proposed method misses the first oscillation count in each weight history.
+9 shows the convergence behavior of the full precision weights and the quantized weights at target-bit.
+We note how the Tiny ViT displays a peculiar convergence behavior, where the accuracy breaks, only to go up again.
+In the Tiny Vit model we regularize the self-attention layers.
+It is already noted in  vitoscillations that ViTs are especially vulnerable to oscillations in the query and key of self-attention layers, which might be related to the convergence behavior observed when regularizing with Eq.
+19.
+A.6 Weight Distribution and Oscillation Frequency
+QAT $\lambda=0$ $\lambda=10$
+Oscillation Weights Oscillation Weights Oscillation Weights
+1_0_conv1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+2_0_conv1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+4_1_conv2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+fc Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 10: Oscillations count and weight distribution side by side across layers and regularization settings.
+Each pair (left: oscillation, right: weights) corresponds to the same $\lambda$ configuration.
+Note that weight distribution plots do not share x and y axis.